### v1

# FedAdamW: A Communication-Efficient Optimizer with Convergence and Generalization Guarantees for Federated Large Models

Abstract
AdamW has become one of the most effective optimizers for training large-scale models. We have also observed its effectiveness in the context of federated learning (FL). However, directly applying AdamW in federated learning settings poses significant challenges: (1) due to data heterogeneity, AdamW often yields high variance in the second-moment estimate $\boldsymbol{v}$; (2) the local overfitting of AdamW may cause client drift; and (3) Reinitializing moment estimates ($\boldsymbol{v}$, $\boldsymbol{m}$) at each round slows down convergence. To address these challenges, we propose the first Federated AdamW algorithm, called FedAdamW, for training and fine-tuning various large models. FedAdamW aligns local updates with the global update using both a local correction mechanism and decoupled weight decay to mitigate local overfitting. FedAdamW efficiently aggregates the mean of the second-moment estimates to reduce their variance and reinitialize them. Theoretically, we prove that FedAdamW achieves a linear speedup convergence rate of $\mathcal{O}(\sqrt{(L\Delta\sigma_{l}^{2})/(SKR\epsilon^{2})}+(L\Delta)/R)$ without heterogeneity assumption, where $S$ is the number of participating clients per round, $K$ is the number of local iterations, and $R$ is the total number of communication rounds. We also employ PAC-Bayesian generalization analysis to explain the effectiveness of decoupled weight decay in local training. Empirically, we validate the effectiveness of FedAdamW on language and vision Transformer models. Compared to several baselines, FedAdamW significantly reduces communication rounds and improves test accuracy. The code is available in https://github.com/junkangLiu0/FedAdamW.
Introduction
With the rapid growth of data and rising concerns over user privacy, traditional centralized training paradigms have become inadequate. Federated Learning (FL) (McMahan et al. 2017) offers a scalable and privacy-preserving framework that enables collaborative model training across decentralized clients without sharing raw data (Liu et al. 2024, 2025b, 2025a). As data becomes increasingly siloed, FL is a practical solution for large-scale distributed deep learning.
However, recent trends in model design-particularly the rise of large-scale architectures such as GPT (Radford et al. 2018), RoBERTa (Liu et al. 2019), and Vision Transformers (ViT) (Dosovitskiy et al. 2020)—pose new challenges for existing FL algorithms. Specifically, the widely-used FedAvg algorithm, which relies on stochastic gradient descent (SGD(Bottou 2010) in local, struggles to efficiently train Transformer models. This is due to the slow convergence and poor adaptivity of SGD in Transformer models (Zhang et al. 2024a; Liu et al. 2025a), which have more complex architectures compared to CNNs. For example, components such as query, key, and value often require different learning rates to be trained effectively (Zhang et al. 2024a). In contrast, AdamW (Loshchilov, Hutter et al. 2017), an adaptive optimizer with decoupled weight decay, has demonstrated superior performance in centralized training of large models based on Transformer (Vaswani et al. 2017; Liu et al. 2019), offering faster convergence and improved generalization, compared to Adam (Kingma and Ba 2014) and SGD.
Empirically, we also observe this advantage in FL: as shown in Figure 1, local training with AdamW (Local AdamW) converges significantly faster than Local SGD  (McMahan et al. 2017) for training various Transformer models . However, naively applying AdamW in FL leads to the following new challenges:
  • Challenge 1: High variance in second-moment estimate ($\boldsymbol{v}$). Due to non-i.i.d. data across clients, gradient noise leads to high variance in second-moment estimate.
  • Challenge 2: Local overfitting and client drift. While AdamW accelerates local training, it intensifies local overfitting. Under non-i.i.d. data, this manifests as client drift, severely hindering the global model’s performance.
  • Challenge 3: Moment estimate reinitialization. reinitializing the first- and second-moment estimates from scratch in every round hinders the convergence rate.
(a) ViT on CIFAR100
(b) GPT2 on Shakespeare
(c) BERT on CoLA
Figure 1: Performance of Local SGD and Local AdamW. For training ViT-Base, GPT2, and BERT (Liu et al. 2019), we carefully tune the learning rate. For training all these Transformer models, Local SGD is still significantly worse than Local AdamW.
These challenges motivate us to develop Federated AdamW (FedAdamW), a novel optimizer tailored for federated learning. FedAdamW addresses the above issues through two key designs: (1) a local correction mechanism that integrates global gradient estimates into the local update, effectively aligning local and global updates to reduce client drift; and (2) a moment aggregation strategy that aggregates the mean of second-moment estimates is theoretically grounded in the Hessian block structure, to reduce variance of $\boldsymbol{v}$ and avoid repeated initialization.
Empirical results on ViT and LLMs confirm that FedAdamW improves test accuracy and reduces communication overhead compared to strong FL baselines.
Our contributions are summarized as follows:
  • Empirical importance of AdamW and challenges in FL. We empirically demonstrate the effectiveness of the AdamW optimizer in federated settings, particularly for training Transformer models. Our analysis reveals three key challenges when applying AdamW in FL: (1) high variance in second-moment estimation, (2) local overfitting leading to client drift, and (3) inefficient convergence due to repeated initialization of optimizer states.
  • We propose FedAdamW, a principled FL algorithm tailored for adaptive optimizers. To address the above challenges, FedAdamW integrates global update estimate into local updates to mitigate overfitting and improve consistency. Inspired by the Hessian structure, we design a communication-efficient aggregation strategy that communicates the mean of second-moment estimates across clients.
  • Theoretical guarantees with improved convergence and generalization. FedAdamW achieves a linear speedup convergence rate of $\mathcal{O}(\sqrt{(L\Delta\sigma_{l}^{2})/(SKR\epsilon^{2})}+(L\Delta)/R)$. To the best of our knowledge, this is the first federated adaptive optimization algorithm without requiring gradient heterogeneity assumption. Furthermore, we utilize the PAC-Bayesian theory to provide insights into the generalization benefits of decoupled weight decay and global-local alignment.
Related Work
$\bullet$ Heterogeneity Issues in Federated Learning. Data heterogeneity across clients is a fundamental challenge in FL. A range of algorithms have been proposed to mitigate the adverse effects of non-i.i.d. data distributions. For example, FedProx (Li et al. 2020) introduces a proximal term to restrict local updates; SCAFFOLD (Karimireddy et al. 2020) applies control variates to correct client drift; and FedCM (Xu et al. 2021) leverages client momentum to stabilize updates. These methods are predominantly built upon the SGD optimizer, and thus inherit its limitations when optimizing large-scale or deep Transformer-based models.
$\bullet$ Adaptive Optimization in Centralized Settings. Adaptive gradient methods have demonstrated superior empirical performance over SGD in centralized settings, particularly for deep neural networks. Pioneering works include Adagrad (Duchi, Hazan, and Singer 2011), Adadelta (Zeiler 2012), Adam (Kingma and Ba 2014), AMSGrad (Reddi, Kale, and Kumar 2019), and AdamW (Loshchilov, Hutter et al. 2017). AdamW, in particular, decouples weight decay from gradient updates, offering improved generalization and training stability—attributes especially critical for Transformer models (Liu et al. 2019; Zhang et al. 2024a).
$\bullet$ Adaptive Optimization in Federated Learning. Recent efforts have explored integrating adaptive methods into FL. FedOpt (Reddi et al. 2020) incorporates server-side adaptivity using Adam and Yogi. FAFED (Wu et al. 2023) aggregates both the first- and second-moment estimates of Adam across clients to stabilize training. FedAMS  (Chen, Li, and Li 2020) shows that averaging the second-moment estimate of Adam is crucial to prevent divergence. More recently, Sun et al. (2023) proposed to only aggregate the second-moment estimate to reduce communication overhead. However, these works only conducted experiments on CNN models. These studies are all based on Adam, which performs poorly with large weight decay.
$\bullet$ Our Contribution for Large Models. We propose FedAdamW, the first federated learning algorithm based on AdamW tailored for Transformer models. Our method integrates a global-local update alignment strategy and a communication-efficient aggregation of second-moment estimates. Furthermore, we provide theoretical convergence and PAC-Bayesian generalization analysis, offering new insights into adaptive optimization in FL.
Algorithm 1 Local AdamW Algorithm
1:Initial model $\boldsymbol{x}^{0}$, $\beta_{1}=0.9,\beta_{2}=0.999,\epsilon=10^{-8}$, time step $t\leftarrow 0$, the number of all clients NN, each round selected clients $S$, weight decay $\lambda$.
2:for $r=1,\dots,R$ do
3:  for each selected client $i\in\{1,\dots,S\}$ in parallel do
4:   $\boldsymbol{x}_{i}^{r,0}\leftarrow\boldsymbol{x}^{r}$, $\boldsymbol{m}^{r,0}_{i}\leftarrow\boldsymbol{0}$, $\boldsymbol{v}^{r,0}_{i}\leftarrow\boldsymbol{0}$;
5:   for $k=1,\dots,K$ do
6:    $\boldsymbol{g}^{r,\tau}_{i}\leftarrow\nabla f_{i}(\boldsymbol{x}_{i}^{r,k};\xi_{i})$;
7:    $\boldsymbol{m}^{r,k}_{i}=\beta_{1}\boldsymbol{m}^{r,k-1}_{i}+\left(1-\beta_{1}\right)\boldsymbol{g}^{r,k}_{i}$;
8:    $\boldsymbol{v}^{r,k}_{i}=\beta_{2}\boldsymbol{v}^{r,k-1}_{i}+\left(1-\beta_{2}\right)\boldsymbol{g}^{r,k}_{i}\odot\boldsymbol{g}^{r,k}_{i}$;
9:    $\hat{\boldsymbol{m}}^{r,k}_{i}=\boldsymbol{m}^{r,k}_{i}/\left(1-\beta_{1}^{k}\right)$;
10:    $\hat{\boldsymbol{v}}^{r,k}_{i}=\boldsymbol{v}^{r,k}_{i}/\left(1-\beta_{2}^{k}\right)$;
11:    $\boldsymbol{x}_{i}^{r,k+1}\!=\!\boldsymbol{x}_{i}^{r,k}-\eta(\hat{\boldsymbol{m}}^{r,k}_{i}/(\sqrt{\hat{\boldsymbol{v}}^{r,k}_{i}}+\epsilon)-\lambda\boldsymbol{x}_{i}^{r,k})$;
12:   end for
13:   Communicate $(\boldsymbol{x}^{r,K}_{i}-\boldsymbol{x}^{r,0}_{i})$ to Server;
14:  end for
15:  $\boldsymbol{x}^{r+1}=\boldsymbol{x}^{r}+\frac{1}{S}\sum_{i=1}^{S}(\boldsymbol{x}^{r,K}_{i}-\boldsymbol{x}^{r,0}_{i})$;
16:  Communicate $(\boldsymbol{x}^{r+1})$ to Clients.
17:end for
Algorithm 2 FedAdamW Algorithm
1:Initial model $\boldsymbol{x}^{0}$, $\beta_{1}=0.9,\beta_{2}=0.999,\epsilon=10^{-8}$, time step $t\leftarrow 0$, the number of all clients NN, each round selected clients $S$, weight decay $\lambda$.
2:for $r=1,\dots,R$ do
3:  for each selected client $i\in\{1,\dots,S\}$ in parallel do
4:   $\boldsymbol{x}_{i}^{r,0}\leftarrow\boldsymbol{x}^{r}$, $\boldsymbol{m}^{r,0}_{i}\leftarrow\boldsymbol{0}$, $\boldsymbol{v}^{r,0}_{i}\leftarrow\boldsymbol{\bar{v}}^{r}$;
5:   for $k=1,\dots,K$ do
6:    $t\leftarrow t+1$;
7:    $\boldsymbol{g}^{r,k}_{i}\leftarrow\nabla f_{i}(\boldsymbol{x}_{i}^{r,k};\xi_{i})$;
8:    $\boldsymbol{m}^{r,k}_{i}=\beta_{1}\boldsymbol{m}^{r,k-1}_{i}+\left(1-\beta_{1}\right)\boldsymbol{g}^{r,k}_{i}$;
9:    $\boldsymbol{v}^{r,k}_{i}=\beta_{2}\boldsymbol{v}^{r,k-1}_{i}+\left(1-\beta_{2}\right)\boldsymbol{g}^{r,k}_{i}\odot\boldsymbol{g}^{r,k}_{i}$;
10:    Bias correction
11:    $\hat{\boldsymbol{m}}^{r,k}_{i}=\boldsymbol{m}^{r,k}_{i}/\left(1-\beta_{1}^{k}\right)$;
12:    $\hat{\boldsymbol{v}}^{r,k}_{i}=\boldsymbol{v}^{r,k}_{i}/\left(1-\beta_{2}^{t}\right)$;
13:    $\vartheta_{i}^{r,k}=1/(\sqrt{\hat{\boldsymbol{v}}_{i}^{r,k}}+\epsilon)$;
14:    Update model parameters $\boldsymbol{x}_{i}^{r,k+1}=\boldsymbol{x}_{i}^{r,k}-\eta(\hat{\boldsymbol{m}}^{r,k}_{i}\odot\vartheta_{i}^{r,k}+\alpha\boldsymbol{\Delta}_{G}^{r}-\lambda\boldsymbol{x}_{i}^{r,k})$;
15:   end for
16:   Communicate $(\boldsymbol{x}^{r,K}_{i}-\boldsymbol{x}^{r,0}_{i},\boldsymbol{\bar{v}}_{i}=\texttt{mean}\Big(\boldsymbol{v}^{r,K}_{i}\Big))$ to Server;
17:  end for
18:  $\boldsymbol{\Delta}_{G}^{r}=\frac{-1}{SK\eta}\sum_{i=1}^{S}(\boldsymbol{x}^{r,K}_{i}-\boldsymbol{x}^{r,0}_{i})$;
19:  $\boldsymbol{x}^{r+1}=\boldsymbol{x}^{r}+\frac{1}{S}\sum_{i=1}^{S}(\boldsymbol{x}^{r,K}_{i}-\boldsymbol{x}^{r,0}_{i})$;
20:  $\boldsymbol{\bar{v}}^{r+1}=\frac{1}{S}\sum_{i=1}^{S}\boldsymbol{\bar{v}}_{i}$;
21:  Communicate $(\boldsymbol{x}^{r+1},\boldsymbol{\bar{v}}^{r+1},\boldsymbol{\Delta}_{G}^{r})$ to Clients.
22:end for
FL Problem Setup
FL aims to optimize model parameters with local clients, i.e., minimizing the following population risk:
$f(\boldsymbol{\boldsymbol{x}})=\frac{1}{N}\sum_{i=1}^{N}\left(f_{i}(\boldsymbol{\boldsymbol{x}}):=\mathbb{E}_{\xi_{i}\sim\mathcal{D}_{i}}\left[F_{i}\left(\boldsymbol{\boldsymbol{x}};\xi_{i}\right)\right]\right).$ (1)
The function $f_{i}$ represents the loss function on client $i$. $\mathbb{E}_{\xi_{i}\sim\mathcal{D}_{i}}[\cdot]$ denotes the conditional expectation with respect to the sample $\xi_{i}$. $\xi_{i}$ is drawn from distribution $\mathcal{D}_{i}$ in client $i$. NN is the number of clients.
(a) High Variance in $\boldsymbol{v}$
(b) Local AdamW client drift
Figure 2: Training on CIFAR-100 using ViT-Tiny. (a) Data heterogeneity causes high variance in second-moment estimates across clients of Local AdamW. (b) Local AdamW suffers from more severe client drift than Local SGD under non-i.i.d. data.
Challenges of AdamW in FL
Despite the widespread use of AdamW (Loshchilov, Hutter et al. 2017; Vaswani et al. 2017) in centralized deep learning, its adaptation to federated settings remains largely unexplored. In this section, we analyze three fundamental challenges that hinder its effectiveness in FL settings.
Challenge 1: High Variance in Second-Moment Estimates ($\boldsymbol{v}$). AdamW maintains a second-moment estimate ($\boldsymbol{v}$) to scale gradients adaptively, updated as:
$\boldsymbol{v}^{r,k}_{i}=\beta_{2}\boldsymbol{v}^{r,k-1}_{i}+(1-\beta_{2})\boldsymbol{g}^{r,k}_{i}\odot\boldsymbol{g}^{r,k}_{i},$ (2)
where $\boldsymbol{v}^{r,k}_{i}$ denotes the second-moment estimate maintained by client $i$ at local step $k$ of round $r$, $\boldsymbol{g}^{r,k}_{i}$ is the stochastic gradient, $\beta_{2}=0.999$ is the exponential decay rate for the second moment, and $\odot$ represents the element-wise (Hadamard) product in Algorithm 1. In FL, data heterogeneity leads to gradient heterogeneity. The squared stochastic gradients $\boldsymbol{g}^{r,k}_{i}\odot\boldsymbol{g}^{r,k}_{i}$ in Eq. (2) amplify the variance of $\boldsymbol{v}$ across clients in Figure 2 (a). This can cause instability and inefficient aggregation, especially when using non-i.i.d. data (Chen, Li, and Li 2020).
Challenge 2: Local Overfitting and Client Drift. While AdamW accelerates convergence through its adaptivity, it may exacerbate local overfitting. In FL, where each client minimizes its own local objective $f_{i}(\cdot)$, creating a natural gap between the local and global optima. Adaptive optimizers such as AdamW, with stronger update magnitudes, can drive clients further toward their local optima—diverging from the global direction. This leads to client drift as illustrated in Figure 2 (b), which manifests as inconsistencies in local models that degrade the global aggregation performance.
Challenge 3: Reinitialization Overhead. In FL, AdamW optimizer states are reinitialized from zero each round:
$\boldsymbol{m}^{r,0}_{i}\leftarrow\boldsymbol{0},\quad\boldsymbol{v}^{r,0}_{i}\leftarrow\boldsymbol{0}.$ (3)
Reinitializing moment estimates across rounds erases temporal memory, hindering the accumulation of adaptive statistics and slowing convergence, particularly in deep or large-scale models.
Our Algorithm: FedAdamW
(a) query (4 heads)
(b) key (4 heads)
(c) value (4 heads)
Figure 3: (a–c):Block-wise Hessian structure of Transformer parameters under FL. Visualizing the Hessian submatrices of query, key, and value heads. The near block-diagonal structure supports block-wise second-moment aggregation in FedAdamW.
Figure 4: Illustration of FedAdamW’s block-wise aggregation strategy Clients estimate local second-moment statistics and send block-wise means to the server, reducing communication cost.
Based on theoretical motivation, we propose an efficient improvement to AdamW called Federated AdamW (FedAdamW). To address Challenge 1, it was experimentally discovered that aggregating AdamW second-moment estimate can stabilize the training process (see Table 7 below). However, aggregating second-moment estimate leads to a double communication.
(Q1) How to efficiently aggregate $\boldsymbol{v}$?
We observe that the Hessian matrix in neural networks exhibits an approximate block-diagonal structure with several dense sub-blocks (Collobert 2004; Zhang et al. 2024b) as shown in Figure 3. In such a structure, a single learning rate can effectively capture the curvature within each block. Leveraging this, we propose a communication-efficient strategy that partitions the second-moment estimate $\boldsymbol{v}$ into $B$ blocks and transmits only the mean of each in Figure 4:
$\boldsymbol{\bar{v}}_{b}=\texttt{mean}(\boldsymbol{v}_{b}),\quad b=1,\dots,B.$ (4)
Block-wise Partitioning Strategy (ViT Example).
We group the parameters into semantically aligned classes that exhibit similar curvature patterns:
  • Class 1: query and key. Query and Key parameters. Each block corresponds to one attention head, as shown in Figure 3.
  • Class 2: attn.proj and MLPs. Blocks align with output neurons in projection and feedforward layers.
  • Class 3: value. Structure is less regular but still shows diagonal blocks; curvature magnitude is notably higher (up to $10^{6}\times$), possibly due to its position after softmax.
  • Class 4: Embedding and output layers. Sub-blocks align with input tokens, forming near-diagonal Hessians.
CNNs (e.g., ResNet): Blocked by convolutional layers or residual blocks. This reduces the communication cost from billions of scalars to $B$ values while preserving adaptive behavior. Empirically, we find this approach also improves generalization in local optimization as shown in Table 7 below. See Appendix D for block partitioning details.
(Q2) How to overcome overfitting in Local AdamW?
To address local overfitting (i.e., Challenge 2), we adopt a stronger weight decay. Unlike Adam, AdamW employs decoupled weight decay, which improves generalization, particularly in federated settings (see Table 6 below). To further mitigate client drift under non-i.i.d. data, we incorporate a global update estimate into the local update rule:
$\boldsymbol{x}_{i}^{r,k+1}=\boldsymbol{x}_{i}^{r,k}-\eta\left(\hat{\boldsymbol{m}}^{r,k}_{i}\odot\vartheta_{i}^{r,k}-\lambda\boldsymbol{x}_{i}^{r,k}+\alpha\boldsymbol{\Delta}_{G}^{r}\right),$ (5)
where $\boldsymbol{\Delta}_{G}^{r}\!=\!\frac{-1}{SK\eta}\sum_{i=1}^{S}(\boldsymbol{x}_{i}^{r,K}\!-\!\boldsymbol{x}_{i}^{r,0})$ is the estimated global update. As shown in Figure 5, this alignment reduces the divergence of local models and improves global consistency.
(Q3) How to initialize second-moment estimates?
We find that initializing the second-moment estimate $\boldsymbol{v}$ with its aggregated mean $\boldsymbol{v}$ significantly accelerates convergence (see Table 7 below). In contrast, we reinitialize the first-moment estimate $\boldsymbol{m}$ to zero at each round. This is because $\boldsymbol{m}$ adapts quickly to recent gradients and does not require long-term accumulation to remain effective.
Figure 5: An illustration of local update in FedAdamW, which corrects client drift caused through global update guidance.
Theoretical Analysis
Convergence Analysis
In this part, we give the convergence theoretical analysis of our proposed FedAdamW algorithm. Firstly we state some standard assumptions for the non-convex function $f$.
Assumption 1 (Smoothness).
(Smoothness) The non-convex $f_{i}$ is a $L$-smooth function for all $i\in[m]$, i.e., $\|\nabla f_{i}(\boldsymbol{x})-\nabla f_{i}(\boldsymbol{y})\|\leq L\|\boldsymbol{x}-\boldsymbol{y}\|$, for all $\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{d}$.
Assumption 2 (Bounded Stochastic Gradient).
$\boldsymbol{g}_{i}^{r}=\nabla f_{i}(\boldsymbol{x}_{i}^{r},\xi_{i}^{r})$ computed by using a sampled mini-batch data $\xi_{i}^{r}$ in the local client $i$ is an unbiased estimator of $\nabla f_{i}$ with bounded variance, i.e., $\mathbb{E}_{\xi_{i}^{r}}[\boldsymbol{g}_{i}^{r}]=\nabla f_{i}(\boldsymbol{x}_{i}^{r})$ and $\mathbb{E}_{\xi_{i}^{r}}\|g_{i}^{r}-\nabla f_{i}(\boldsymbol{x}_{i}^{r})\|^{2}\leq\sigma_{l}^{2}$, for all $\boldsymbol{x}_{i}^{r}\in\mathbb{R}^{d}$.
Assumption 3 (Bounded Stochastic Gradient II).
Each element of stochastic gradient $\boldsymbol{g}_{i}^{r}$ is bounded, i.e., $\|\boldsymbol{g}_{i}^{r}\|_{\infty}=\|f_{i}(\boldsymbol{x}_{i}^{r},\xi_{i}^{r})\|_{\infty}\leq G_{g}$, for all $\boldsymbol{x}_{i}^{r}\in\mathbb{R}^{d}$ and any sampled mini-batch data $\xi_{i}^{r}$.
Assumption 4 (Bounded Heterogeneity).
The dissimilarity between local clients is bounded on the gradients, i.e., $\|\nabla f_{i}(\boldsymbol{x})-\nabla f(\boldsymbol{x})\|^{2}\leq\sigma_{g}^{2}$, for all $\boldsymbol{x}\in\mathbb{R}^{d}$.
These assumptions are standard in federated adaptive optimization literature (Fan et al. 2024; Sun et al. 2023).
Theorem 1 (Convergence for non-convex functions).
Under Assumptions 1, 2, and 3, if we take $g^{0}=0$$\beta_{1}=0,\lambda=0$ then FedAdamW converges as follows
$\frac{1}{R}\sum_{r=0}^{R-1}\mathbb{E}\left[\left\|\nabla f\left(\boldsymbol{x}^{r}\right)\right\|^{2}\right]\lesssim\mathcal{O}\left(\sqrt{\frac{L\Delta\sigma_{l}^{2}}{SKR\epsilon^{2}}}+\frac{L\Delta}{R}\right).$ (6)
Here $G_{0}:=\frac{1}{N}\sum_{i=1}^{N}\left\|\nabla f_{i}\left(\boldsymbol{x}^{0}\right)\right\|^{2}$$\Delta=f\left(\boldsymbol{x}^{0}\right)-f^{\star}$, $S$ is the number of participating clients per round, $K$ is the number of local iterations, and $R$ is the total number of communication rounds.
The proof is provided in Appendix A. The convergence rate of FedAdamW is faster than that of Local AdamW and FedLADA’s $\mathcal{O}\left(\sqrt{\frac{L\Delta(\sigma_{l}^{2}+\sigma_{g}^{2})}{SKR\epsilon^{2}}}+\frac{L\Delta}{R}\right)$, and we do not need Assumption 4. This is due to the suppression of local drift by the global update estimation $\boldsymbol{\Delta}_{G}^{r}$. We have verified this in Table 5 below.
Generalization Analysis
Theorem 2.
Assume the prior hypothesis $\boldsymbol{x}_{0}$ satisfies $\mathcal{P}_{\text{pre }}\sim\mathcal{N}(\mathbf{0},\rho\boldsymbol{I})$. Then the expected risk for the posterior hypothesis $\boldsymbol{x}\sim\mathcal{P}$ of FedAdamW learned on training dataset $\mathcal{D}_{tr}\sim\mathcal{D}$ with $n$ samples holds
$\mathbb{E}_{\boldsymbol{\xi}\sim\mathcal{D},\boldsymbol{x}\sim\mathcal{P}}[f(\boldsymbol{x},\boldsymbol{\xi})]-\mathbb{E}_{\boldsymbol{\xi}\in\mathcal{D}_{tr},\boldsymbol{x}\sim\mathcal{P}}[f(\boldsymbol{x},\boldsymbol{\xi})]\leq$
$\frac{\sqrt{8}}{\sqrt{n}}\left(\sum_{i=1}^{d}\log\frac{2\rho b\left(\sigma_{i}^{\frac{1}{2}}+\lambda\right)}{\eta}\!+\!\frac{\eta}{2\rho b}\sum_{i=1}^{d}\frac{1}{\sigma_{i}^{\frac{1}{2}}+\lambda}\!+\!c_{0}\!\right)^{\frac{1}{2}}$
, with at least probability $1-\tau$, where $\tau\in(0,1)$ and $c_{0}=\frac{1}{2\rho}\left\|\boldsymbol{x}_{*}\right\|^{2}-\frac{d}{2}+2\ln\left(\frac{2n}{\tau}\right)$. Here, $\sigma_{i}$ represents the local curvature (e.g., Hessian eigenvalue). $b$ is the batch size, $\eta$ is the learning rate, $\lambda$ is a weight decay parameter, $n$ is the training set size, and $d$ is the parameter dimension.
The proof is provided in Appendix B. Theorem 2 shows that the generalization error of FedAdamW can be upper bounded by $\mathcal{O}(1/\sqrt{n})$, where $n$ is the number of tatal data, consistent with classical results from PAC theory, stability, and uniform convergence (Shalev-Shwartz and Ben-David 2014). We further analyze the impact of the decoupled weight decay parameter $\lambda$ on this bound. As $\lambda$ increases, the first term $\sum_{i=1}^{d}\log 2\rho b(\sigma_{i}^{1/2}+\lambda)\eta^{-1}$ increases, while the second term $\frac{\eta}{2\rho b}\sum_{i=1}^{d}(\sigma_{i}^{1/2}+\lambda)^{-1}$ decreases. Although choosing the optimal $\lambda$ is challenging in practice, this trade-off suggests that tuning $\lambda$ appropriately can lead to a smaller generalization error, as shown in Table Table 6 below. This explains why FedAdamW often outperforms Local Adam (which corresponds to $\lambda=0$).
(a) ResNet18, Dir-0.6
(b) ResNet18, Dir-0.1
(c) ViT-Tiny, Dir-0.6
(d) ViT-Tiny, Dir-0.1
Figure 6: Training loss curves on CIFAR-100 using ResNet-18 and ViT-Tiny under Dir-0.1, Dir-0.6. FedAdamW converges the fastest.
Table 1: Test accuracy, training loss, and communication cost of each method on CIFAR-100 using ResNet-18 and ViT-Tiny over 300 communication rounds under Dir-0.6 and Dir-0.1 settings (100 clients, 10% participation, batch size 50, $K=50$).
Method ResNet-18 (Dir-0.6) ResNet-18 (Dir-0.1) ViT-Tiny (Dir-0.6) ViT-Tiny (Dir-0.1) Comm
Test Acc Train Loss Test Acc Train Loss Test Acc Train Loss Test Acc Train Loss
FedAvg $64.08_{\pm 0.18}$ 0.376 $60.25_{\pm 0.20}$ 0.767 $32.36_{\pm 0.08}$ 2.350 $27.14_{\pm 0.12}$ 2.867 $1\times$
SCAFFOLD $65.01_{\pm 0.15}$ 0.365 $59.37_{\pm 0.16}$ 0.814 $32.17_{\pm 0.12}$ 2.295 $27.31_{\pm 0.11}$ 2.855 $2\times$
FedCM $48.69_{\pm 0.10}$ 1.305 $44.43_{\pm 0.08}$ 1.645 $26.33_{\pm 0.06}$ 2.681 $23.18_{\pm 0.15}$ 3.038 $1\times$
Local Adam $60.98_{\pm 0.28}$ 1.598 $58.88_{\pm 0.20}$ 0.975 $38.69_{\pm 0.16}$ 2.082 $29.88_{\pm 0.08}$ 2.961 $1\times$
FedAdam $63.77_{\pm 0.13}$ 0.562 $61.62_{\pm 0.16}$ 0.707 $28.77_{\pm 0.12}$ 2.709 $23.49_{\pm 0.15}$ 3.084 $1\times$
FedLADA $65.07_{\pm 0.18}$ 0.671 $59.93_{\pm 0.21}$ 0.556 $37.31_{\pm 0.16}$ 2.127 $35.33_{\pm 0.16}$ 2.678 $2\times$
Local AdamW $62.84_{\pm 0.08}$ 0.363 $58.97_{\pm 0.10}$ 0.794 $40.47_{\pm 0.09}$ 1.026 $36.86_{\pm 0.11}$ 1.954 $1\times$
FedAdamW 66.12±0.10 0.122 63.01±0.12 0.480 42.56±0.10 0.401 39.86±0.16 1.251 $1\times$
Experiments
Datasets. We evaluate FedAdamW on both vision and language tasks. (i) For image classification, we use CIFAR-100 (Krizhevsky, Hinton et al. 2009), and Tiny ImageNet (Le and Yang 2015). (ii) For NLP tasks, we adopt benchmark datasets from the GLUE benchmark, including SST-2 (Socher et al. 2013), QQP (Socher et al. 2013). To simulate data heterogeneity across clients, we follow the Dirichlet partitioning scheme (Hsu, Qi, and Brown 2019), where a Dir-0.6 corresponds to a low heterogeneity and Dir-0.1 implies high heterogeneity.
Model Architectures. We explore a variety of model types: (i) ResNet-18 (He et al. 2016) as a representative convolutional neural network (CNN), (ii) Swin Transformer (Liu et al. 2021) and ViT-Tiny (Dosovitskiy et al. 2020) for Vision Transformers, and (iii) RoBERTa-Base (Liu et al. 2019) for large-scale language model.
Baselines. We compare our method against state-of-the-art FL algorithms: FedAvg (McMahan et al. 2017), SCAFFOLD (Karimireddy et al. 2020), FedCM (Xu et al. 2021), FedAdam (Reddi et al. 2020), FedLADA (Sun et al. 2023), Local Adam and Local AdamW.
Hyperparameter Settings. For FedAvg, SCAFFOLD, FedCM, FedAdam, the $lr$ is selected from $\{10^{-2},\ 3\times 10^{-2},\ 5\times 10^{-2},\ 10^{-1},\ 3\times 10^{-1}\}$, with a weight decay of $0.001$. For FedAdamW, FedLADA, Local Adam and Local AdamW, the $lr$ is selected from $\{10^{-4},\ 3\times 10^{-4},\ 5\times 10^{-4},\ 8\times 10^{-4},\ 10^{-3}\}$, with weight decay $0.01$ or $0.001$, $\beta_{1}=0.9$, $\beta_{2}=0.999$. We apply cosine learning rate decay, and set FedAdamW to $\boldsymbol{\alpha}\!=\!0.5$, weight decay $\boldsymbol{\lambda}\!=\!0.01$. Additional hyperparameter configurations are detailed in Appendix C. We release all code, configuration files to ensure full reproducibility. All results are averaged over 5 runs with std reported.
Questions. Our experiments are designed to answer the following: Q1. Does Local AdamW outperform Local SGD when training Transformer models? Q2. Can FedAdamW effectively address the three challenges identified for AdamW in FL? Q3. Is FedAdamW generally effective across both CNNs and Transformers? Q4. Are individual components of FedAdamW—such as global update correction, decoupled weight decay, and block-wise $\boldsymbol{v}$ averaging—empirically beneficial? Q5. Do our theoretical findings (Theorems 1 and 2) align with empirical results?
Table 2: Comparison of test accuracy and training loss for Swin Transformer under Dir-0.1 with 100 communication rounds (100 clients, 5% participation, batch size 16, $K=50$).
Method CIFAR-100 Tiny ImageNet
Test Acc Train Loss Test Acc Train Loss
FedAvg $80.02_{\pm 0.28}$ 0.588 $80.38_{\pm 0.22}$ 0.826
SCAFFOLD $81.30_{\pm 0.18}$ 0.514 $82.41_{\pm 0.18}$ 0.650
FedCM $82.38_{\pm 0.19}$ 0.565 $83.18_{\pm 0.19}$ 0.522
Local Adam $79.75_{\pm 0.26}$ 0.534 $73.63_{\pm 0.28}$ 1.045
FedAdam $77.48_{\pm 0.19}$ 0.651 $78.20_{\pm 0.22}$ 0.834
FedLADA $74.64_{\pm 0.18}$ 0.598 $70.95_{\pm 0.19}$ 0.944
Local AdamW $83.35_{\pm 0.10}$ 0.381 $80.26_{\pm 0.12}$ 0.686
FedAdamW 85.85±0.08 0.285 85.23±0.10 0.446
Table 3: Test accuracy (%) using RoBERTa-Base with LoRA across seven GLUE tasks over 100 communication rounds (mean ± std).
Method (Dir-0.8) CoLA RTE SST-2 QQP MRPC QNLI MNLI Avg Acc.
FedAvg 56.12±0.18 48.72±0.25 93.66±0.10 85.87±0.14 86.00±0.12 90.21±0.09 83.22±0.17 77.68±0.17
SCAFFOLD 57.79±0.21 51.62±0.28 93.15±0.11 84.25±0.15 86.11±0.13 90.32±0.10 83.49±0.18 77.82±0.17
FedCM 56.29±0.16 64.98±0.22 93.25±0.12 83.19±0.17 85.56±0.13 88.13±0.15 78.90±0.19 78.33±0.18
Local Adam 56.08±0.19 62.81±0.23 93.80±0.09 85.07±0.13 84.55±0.14 88.57±0.12 82.62±0.16 79.07±0.15
FedAdam 55.26±0.20 58.12±0.26 93.26±0.10 85.12±0.13 86.11±0.11 89.21±0.09 83.16±0.17 78.32±0.16
FedLADA 50.00±0.24 57.40±0.25 93.57±0.11 85.88±0.14 82.59±0.15 89.76±0.10 82.99±0.16 77.17±0.17
Local AdamW 56.45±0.17 54.15±0.27 93.57±0.09 86.93±0.12 86.27±0.11 90.73±0.08 84.26±0.14 78.91±0.15
FedAdamW (ours) 58.21±0.15 66.48±0.20 94.03±0.08 87.62±0.11 86.76±0.10 90.88±0.07 84.55±0.13 81.79±0.14
Results on Convolutional Neural Networks
Training on CIFAR-100 with ResNet-18.
Table 11 and Figure 6 present the test accuracy and training loss on CIFAR-100 using ResNet-18. FedAdamW achieves the best performance under both Dir-0.6 and Dir-0.1 settings, reaching a top accuracy of 66.12% and 63.01%, respectively. It also attains the lowest training loss (0.122 and 0.480), demonstrating faster and more stable convergence. Compared to other adaptive baselines such as FedAdam and Local AdamW, FedAdamW shows superior generalization under data heterogeneity, confirming its effectiveness in CNNs (Q3).
Results on Transformer Models
Training on CIFAR-100 with ViT-Tiny. Table 11 and Figure 6 show FedAdamW achieves the best performance across both heterogeneity levels, with test accuracies of 42.56% (Dir-0.6) and 38.25% (Dir-0.1), and the lowest training loss (0.401 and 1.251), confirming its efficient convergence (Q5). Compared to Local AdamW, it provides consistent improvements in both accuracy and stability (Q1,Q2, Q3). Moreover, other adaptive baselines such as FedAdam and FedLADA perform significantly worse under high heterogeneity, highlighting the effectiveness of global update correction and decoupled weight decay (Q4). These results validate that FedAdamW is particularly effective for federated vision Transformers under non-i.i.d. conditions. The small dataset CIFAR100 is difficult to support the performance of ViT, resulting in lower accuracy. Therefore, we continued to test on the pretrained model.
Fine-tuning Results on Swin Transformer. Table 2 reports results on Swin Transformer under Dir-0.1. FedAdamW achieves the highest test accuracy on both CIFAR-100 (85.85%) and Tiny ImageNet (85.23%), while also attaining the lowest training loss, reflecting faster convergence and improved generalization. Compared to other adaptive baselines such as FedAdam and FedLADA, FedAdamW consistently outperforms across both datasets, demonstrating its effectiveness in fine-tuning large Transformer models under non-i.i.d. conditions.
Fine-tuning Results on LLMs. Table 3 summarizes results on the GLUE benchmark using RoBERTa-Base with LoRA, 20 clients, 20% participation, batch size 32, $K=50$, rank==16. FedAdamW achieves the highest average accuracy of 81.79%, outperforming strong baselines such as FedAvg (77.68%) and Local AdamW (78.91%). It is particularly strong on challenging tasks like RTE and QQP, exceeding the next best methods by +1.50% and +1.74%, respectively.
Ablation Study
Impact of A1, A2, A3. Table 4 summarizes the effect of removing key components in FedAdamW. We draw the following observations: $\bullet$ Removing second-moment aggregation (A1) significantly degrades performance, indicating that mean($\boldsymbol{v}$) aggregation stabilizes adaptive updates across clients. $\bullet$ Without global gradient alignment (A2), the local models drift apart, resulting in higher train loss and lower generalization. $\bullet$ The use of standard (non-decoupled) weight decay (A3) leads to suboptimal regularization, confirming the necessity of decoupled weight decay for Transformer-based federated training. $\bullet$ The complete FedAdamW (A4) consistently outperforms all ablated versions, validating the effectiveness of our joint design.
Impact of $\alpha$. Table 5 evaluates the effect of the global update alignment parameter $\alpha$ in FedAdamW. As predicted by our convergence analysis (Theorem 1), incorporating global update direction helps suppress client drift and accelerates convergence. We observe that $\alpha=0.5$ yields the best performance, striking a balance between local adaptivity and global consistency, in line with our theoretical insight (Q5).
Table 4: Ablation study of FedAdamW on CIFAR-100 using ViT-Tiny (Dir-0.1, 300 rounds).
Variant Test Acc (%) Train Loss
A1: w/o $\boldsymbol{\bar{v}}$ (no moment agg.) $37.51_{\pm 0.12}$ 1.504
A2: w/o $\boldsymbol{\Delta}_{G}$ (no global align.) $37.42_{\pm 0.14}$ 1.621
A3: w/o decoupled weight decay $38.25_{\pm 0.16}$ 1.356
A4: FedAdamW (Full) $\mathbf{39.86}_{\pm 0.16}$ $\mathbf{1.251}$
Table 5: Impact of $\alpha$ on FedAdamW using ViT-Tiny on CIFAR-100 (Dir-0.1).
$\alpha$ 0.00 0.25 0.50 0.75 1.00
Test Acc (%) 36.86 37.93 39.86 37.47 36.25
Train Loss 1.954 1.586 1.251 1.362 1.491
Impact of weight decay $\lambda$. Table 6 shows that decoupled weight decay, as used in AdamW and FedAdamW, consistently improves test accuracy over standard Adam. FedAdamW generalizes well across all $\lambda$ values, with $\lambda=0.01$ performing best. This aligns with our PAC-Bayesian analysis (Theorem 2), where an appropriate $\lambda$ balances regularization and curvature for better generalization (Q5).
Table 6: Ablation on weight decay $\lambda$ using ViT-Tiny on CIFAR-100 (Dir-0.1). FedAdamW consistently outperforms local baselines across a range of $\lambda$ values.
$\lambda$ 0.0005 0.001 0.005 0.010 0.020
Local Adam 28.86 29.88 18.65 8.56 4.05
Local AdamW 35.82 36.12 36.54 36.86 36.28
FedAdamW 38.26 39.24 39.55 39.86 38.56
Impact of Aggregation Strategy. Table 7 shows that our strategy, Agg-mean-$\boldsymbol{v}$, achieves the best balance between accuracy and communication cost. While Agg-$\boldsymbol{v}$ improves performance by reducing variance, full aggregation (Agg-$\boldsymbol{vm}$) introduces excessive communication with marginal gains. In contrast, Agg-mean-$\boldsymbol{v}$ attains similar benefits with only $\mathcal{O}(B)$ communication—where $B$ is the number of blocks—demonstrating its scalability and effectiveness in stabilizing updates under strict communication constraints.
Table 7: Ablation study of moment aggregation strategies of Local AdamW on CIFAR-100 with ViT-Tiny under Dir-0.1.
Aggregation Strategy Acc Train Loss Comm(↑)
NoAgg 36.86±0.11 1.954 5.7M
Agg-$\boldsymbol{m}$ 37.12±0.13 1.854 11.4M
Agg-$\boldsymbol{v}$ 38.01±0.12 1.652 11.4M
Agg-$\boldsymbol{vm}$ (FullAgg) 38.12±0.12 1.645 17.1M
Agg-mean-$\boldsymbol{v}$ 38.15±0.10 1.601 5.7M
Conclusion
In this work, we proposed a novel federated optimization algorithm (FedAdamW) for training large-scale Transformer models. FedAdamW tackles the key challenges of applying AdamW in federated settings, including high variance in second-moment estimates, local overfitting under non-i.i.d. data, and inefficiencies from frequent reinitialization. It integrates second-moment aggregation, global update correction, and decoupled weight decay. We provided convergence analysis under non-convex and use the PAC Bayesian theory to support its generalization benefits. Extensive experiments on vision and language tasks verified that FedAdamW consistently outperforms strong FL baselines, especially on Transformer architectures, demonstrating its practical and theoretical strengths. We believe FedAdamW opens a new direction for adapting modern optimizers to FL such as LAMB (Chen et al. 2023) or Lion (Chen et al. 2023).
Appendix A Appendix A: Proof of Theorem 1 and Convergence Analysis
Algorithm 3 FedAdamW Algorithm with $\lambda=0,\beta_{1}=0$.
1:Initial model $\boldsymbol{x}^{0}$, $\beta_{1}=0,\beta_{2}=0.999,\epsilon=10^{-8}$, time step $t\leftarrow 0$, the number of all clients NN, each round selected clients $S$, weight decay $\lambda$.
2:for $r=1,\dots,R$ do
3:  for each selected client $i\in\{1,\dots,S\}$ in parallel do
4:   $\boldsymbol{x}_{i}^{r,0}\leftarrow\boldsymbol{x}^{r}$, $\boldsymbol{m}^{r,0}_{i}\leftarrow\boldsymbol{0}$, $\boldsymbol{v}^{r,0}_{i}\leftarrow\boldsymbol{\bar{v}}^{r}$;
5:   for $k=1,\dots,K$ do
6:    $t\leftarrow t+1$;
7:    $\boldsymbol{g}^{r,k}_{i}\leftarrow\nabla f_{i}\left(\boldsymbol{x}_{i}^{r,k};\xi_{i}\right)$;
8:    $\boldsymbol{m}^{r,k}_{i}=\boldsymbol{g}^{r,k}_{i}$;
9:    $\boldsymbol{v}^{r,k}_{i}=\beta_{2}\boldsymbol{v}^{r,k-1}_{i}+\left(1-\beta_{2}\right)\boldsymbol{g}^{r,k}_{i}\odot\boldsymbol{g}^{r,k}_{i}$;
10:    Bias correction
11:    $\hat{\boldsymbol{v}}^{r,k}_{i}=\boldsymbol{v}^{r,k}_{i}/\left(1-\beta_{2}^{t}\right)$;
12:    $\vartheta_{i}^{r,k}=1/\Big(\sqrt{\hat{\boldsymbol{v}}_{i}^{r,k}}+\epsilon\Big)$;
13:    Update model parameters $\boldsymbol{x}_{i}^{r,k+1}=\boldsymbol{x}_{i}^{r,k}-\eta\Big(\alpha\boldsymbol{g}^{r,k}_{i}\odot\vartheta_{i}^{r,k}+(1-\alpha)\boldsymbol{\Delta}^{r}_{G}\Big)$;
14:   end for
15:   Communicate $\Big(\boldsymbol{x}^{r,K}_{i}-\boldsymbol{x}^{r,0}_{i},\boldsymbol{\bar{v}}_{i}=\texttt{mean}\Big(\boldsymbol{v}^{r,K}_{i}\Big)\Big)$ to Server;
16:  end for
17:  $\boldsymbol{\Delta}_{G}=\frac{-1}{SK\eta}\sum_{i=1}^{S}(\boldsymbol{x}^{r,K}_{i}-\boldsymbol{x}^{r,0}_{i})$;
18:  $\boldsymbol{x}^{r+1}=\boldsymbol{x}^{r}-\gamma\boldsymbol{\Delta}^{r}_{G}$;
19:  $\boldsymbol{\bar{v}}^{r+1}=\frac{1}{S}\sum_{i=1}^{S}\boldsymbol{\bar{v}}_{i}$;
20:  Communicate $\Big(\boldsymbol{x}^{r+1},\boldsymbol{\bar{v}}^{r+1},\boldsymbol{\Delta}^{r}_{G}\Big)$ to Clients.
21:end for
To simplify the analysis, we consider the iterative rules as in Algorithm 3, where we let $\beta_{1}=0$, $\lambda=0$. The local update takes the following rule:
$\boldsymbol{x}_{i}^{r,k+1}=\boldsymbol{x}_{i}^{r,k}-\eta\Big(\alpha\boldsymbol{g}^{r,k}_{i}\odot\vartheta_{i}^{r,k}+(1-\alpha)\boldsymbol{\Delta}_{G}\Big).$
Algorithm 2 is the algorithm that performs better in practical situations, while Algorithm 3 is the algorithm used for theoretical analysis. Algorithm 2 is the algorithm that performs better in practical situations.
Assumption A.1 (Smoothness).
The non-convex $f_{i}$ is one $L$-smooth function for all $i\in[m]$, i.e., $\|\nabla f_{i}(\boldsymbol{x})-\nabla f_{i}(\boldsymbol{y})\|\leq L\|\boldsymbol{x}-\boldsymbol{y}\|$, for all $\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{d}$.
Assumption A.2 (Bounded Stochastic Gradient I).
$\boldsymbol{g}_{i}^{r}=\nabla f_{i}(\boldsymbol{x}_{i}^{r},\xi_{i}^{r})$ computed by using a sampled mini-batch $\xi_{i}^{r}$ in client $i$ is an unbiased estimator of $\nabla f_{i}$ with bounded variance: $\mathbb{E}_{\xi_{i}^{r}}[\boldsymbol{g}_{i}^{r}]=\nabla f_{i}(\boldsymbol{x}_{i}^{r})$ and $\mathbb{E}_{\xi_{i}^{r}}\|g_{i}^{r}-\nabla f_{i}(\boldsymbol{x}_{i}^{r})\|^{2}\leq\sigma_{l}^{2}$.
Assumption A.3 (Bounded Stochastic Gradient II).
Each element of stochastic gradient $\boldsymbol{g}_{i}^{r}$ is bounded, i.e., $\|\boldsymbol{g}_{i}^{r}\|_{\infty}=\|f_{i}(\boldsymbol{x}_{i}^{r},\xi_{i}^{r})\|_{\infty}\leq G_{g}$, for all $\boldsymbol{x}_{i}^{r}\in\mathbb{R}^{d}$ and any mini-batch $\xi_{i}^{r}$.
Assumption A.4 (Bounded Heterogeneity).
The gradient dissimilarity between clients is bounded: $\|\nabla f_{i}(\boldsymbol{x})-\nabla f(\boldsymbol{x})\|^{2}\leq\sigma_{g}^{2}$, for all $\boldsymbol{x}\in\mathbb{R}^{d}$.
In this section, we give the theoretical analysis of our proposed FedAdamW algorithm. Firstly we state some standard assumptions for the non-convex function $F$.
We use the common gradient bound condition in our proof. Then we can upper bound $G_{\vartheta}$ as:
$G_{\vartheta}\$$=\left\{\|\vartheta_{i}^{r,k}\|_{\infty}\right\}=\left\{\|\frac{1}{\sqrt{(1-\beta_{1}^{k})^{2}/(1-\beta_{2}^{t})\boldsymbol{v}_{i}^{r,k}}+\epsilon}\|_{\infty}\right\}$
$=\left\{\|\frac{1}{\sqrt{(1-\beta_{1}^{k})^{2}/(1-\beta_{2}^{t})(\beta_{2}\boldsymbol{v}^{r,k-1}_{i}+\left(1-\beta_{2}\right)\boldsymbol{g}^{r,k}_{i}\cdot\boldsymbol{g}^{r,k}_{i}})+\epsilon}\|_{\infty}\right\}.$
$\sqrt{(1-\beta_{1}^{k})^{2}/(1-\beta_{2}^{t})(\beta_{2}\boldsymbol{v}^{r,k-1}_{i}+\left(1-\beta_{2}\right)\boldsymbol{g}^{r,k}_{i}\cdot\boldsymbol{g}^{r,k}_{i}}$ is bounded as:
$\|\sqrt{(1-\beta_{1}^{k})^{2}/(1-\beta_{2}^{t})(\beta_{2}\boldsymbol{v}^{r,k-1}_{i}+\left(1-\beta_{2}\right)\boldsymbol{g}^{r,k}_{i}\cdot\boldsymbol{g}^{r,k}_{i}}\|_{\infty}\leq G_{g}.$ (7)
Thus we bound $G_{\vartheta}$ as $\frac{1}{G_{g}}\leq G_{\vartheta}\leq\frac{1}{\epsilon}$.
Main Lemmas
Lemma 1.
[…]
$\mathbb{E}\left[\left\|\sum_{i=1}^{\tau}X_{i}\right\|^{2}\right]\leq\left\|\sum_{i=1}^{\tau}\mu_{i}\right\|^{2}+\tau^{2}\sigma^{2}.$
If they are correlated in the Markov way such that $\mathbb{E}\left[X_{i}\mid X_{i-1},\cdots X_{1}\right]=\mu_{i}$ and $\mathbb{E}\left[\left\|X_{i}-\mu_{i}\right\|^{2}\mid\right$. $\left.\mu_{i}\right]\leq\sigma^{2}$, i.e., the variables $\left\{X_{i}-\mu_{i}\right\}$ form a martingale. Then the following tighter bound holds:
$\mathbb{E}\left[\left\|\sum_{i=1}^{\tau}X_{i}\right\|^{2}\right]\leq 2\mathbb{E}\left[\left\|\sum_{i=1}^{\tau}\mu_{i}\right\|^{2}\right]+2\tau\sigma^{2}.$
Lemma 2.
Given vectors $v_{1},\cdots,v_{N}\in\mathbb{R}^{d}$ and $\bar{v}=\frac{1}{N}\sum_{i=1}^{N}v_{i}$, if we sample $\mathcal{S}\subset\{1,\cdots,N\}$ uniformly randomly such that $|\mathcal{S}|=S$, then it holds that
$\mathbb{E}\left[\left\|\frac{1}{S}\sum_{i\in\mathcal{S}}v_{i}\right\|^{2}\right]=\|\bar{v}\|^{2}+\frac{N-S}{S(N-1)}\frac{1}{N}\sum_{i=1}^{N}\left\|v_{i}-\bar{v}\right\|^{2}.$
Proof.
Letting $\mathbb{I}\{i\in\mathcal{S}\}$ be the indicator for the event $i\in\mathcal{S}_{r}$, we prove this lemma by direct calculation as follows:
$\mathbb{E}\left[\left\|\frac{1}{S}\sum_{i\in\mathcal{S}}v_{i}\right\|^{2}\right]$ $=\mathbb{E}\left[\left\|\frac{1}{S}\sum_{i=1}^{N}v_{i}\mathbb{I}\{i\in\mathcal{S}\}\right\|^{2}\right]$
$=\frac{1}{S^{2}}\mathbb{E}\left[\left(\sum_{i}\left\|v_{i}\right\|^{2}\mathbb{I}\{i\in\mathcal{S}\}+2\sum_{i<j}v_{i}^{\top}v_{j}\mathbb{I}\{i,j\in\mathcal{S}\}\right)\right]$
$=\frac{1}{SN}\sum_{i=1}^{N}\left\|v_{i}\right\|^{2}+\frac{1}{S^{2}}\frac{S(S-1)}{N(N-1)}2\sum_{i<j}v_{i}^{\top}v_{j}$
$=\frac{1}{SN}\sum_{i=1}^{N}\left\|v_{i}\right\|^{2}+\frac{1}{S^{2}}\frac{S(S-1)}{N(N-1)}\left(\left\|\sum_{i=1}^{N}v_{i}\right\|^{2}-\sum_{i=1}^{N}\left\|v_{i}\right\|^{2}\right)$
$=\frac{N-S}{S(N-1)}\frac{1}{N}\sum_{i=1}^{N}\left\|v_{i}\right\|^{2}+\frac{N(S-1)}{S(N-1)}\|\bar{v}\|^{2}$
$=\frac{N-S}{S(N-1)}\frac{1}{N}\sum_{i=1}^{N}\left\|v_{i}-\bar{v}\right\|^{2}+\|\bar{v}\|^{2}.$
Basic Assumptions and Notations
Let $\mathcal{F}^{0}=\emptyset$ and $\mathcal{F}_{i}^{r,k}:=\sigma\left(\left\{x_{i}^{r,j}\right\}_{0\leq j\leq k}\cup\mathcal{F}^{r}\right)$ and $\mathcal{F}^{r+1}:=\sigma\left(\cup_{i}\mathcal{F}_{i}^{r,K}\right)$ for all $r\geq 0$ where $\sigma(\cdot)$ indicates the $\sigma$-algebra. Let $\mathbb{E}_{r}[\cdot]:=\overline{\mathbb{E}}\left[\cdot\mid\mathcal{F}^{r}\right]$ be the expectation, conditioned on the filtration $\mathcal{F}^{r}$, with respect to the random variables $\left\{\mathcal{S}^{r},\left\{\xi_{i}^{r,k}\right\}_{1\leq i\leq N,0\leq k<K}\right\}$ in the $r$-th iteration. We also use $\mathbb{E}[\cdot]$ to denote the global expectation over all randomness in algorithms. Through out the proofs, we use $\sum_{i}$ to represent the sum over $i\in\{1,\ldots,N\}$, while $\sum_{i\in\mathcal{S}^{r}}$ denotes the sum over $i\in\mathcal{S}^{r}$. Similarly, we use $\sum_{k}$ to represent the sum of $k\in\{0,\ldots,K-1\}$. For all $r\geq 0$, we define the following auxiliary variables to facilitate proofs:
$\mathcal{E}_{r}$ $=\mathbb{E}\left[\left\|\nabla f\left(\boldsymbol{x}^{r}\right)\odot\frac{1}{SK}\sum_{i\in S^{r}}\sum_{k=1}^{K}\vartheta_{i}^{r,k}-{\Delta}^{r+1}_{G}\right\|^{2}\right],$
$U_{r}$ $=\frac{1}{NK}\sum_{i}\sum_{k}\mathbb{E}\left[\left\|\boldsymbol{x}_{i}^{r,k}-\boldsymbol{x}^{r}\right\|\right]^{2},$
$\xi_{i}^{r,k}$ $=\mathbb{E}\left[\boldsymbol{x}_{i}^{r,k+1}-\boldsymbol{x}_{i}^{r,k}\mid\mathcal{F}_{i}^{r,k}\right],$
$\Xi_{r}$ $=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\left\|\xi_{i}^{r,0}\right\|^{2}\right].$
Throughout the Appendix, we let $\Delta:=f\left(\boldsymbol{x}^{0}\right)-f^{\star},G_{0}:=\frac{1}{N}\sum_{i}\left\|\nabla f_{i}\left(\boldsymbol{x}^{0}\right)\right\|^{2},\boldsymbol{x}^{-1}:=\boldsymbol{x}^{0}$ and $\mathcal{E}_{-1}:=$ $\mathbb{E}\left[\left\|\nabla f\left(\boldsymbol{x}^{0}\right)-g^{0}\right\|^{2}\right]$. We will use the following foundational lemma for all our algorithms.
Proof of Theorem 1
Lemma 3.
Under Assumption 1 , if $\gamma L\leq\frac{1}{24}$, the following holds all $r\geq 0$ :
$\mathbb{E}\left[f\left(\boldsymbol{x}^{r+1}\right)\right]\leq\mathbb{E}\left[f\left(\boldsymbol{x}^{r}\right)\right]-\frac{11\gamma}{24}\mathbb{E}\left[\left\|\nabla f\left(\boldsymbol{x}^{r}\right)\right\|^{2}\right]+\frac{13\gamma}{24}\mathcal{E}_{r}.$
Proof.
Since $f$ is $L$-smooth, we have
$f\left(\boldsymbol{x}^{r+1}\right)$ $\leq f\left(\boldsymbol{x}^{r}\right)+\left\langle\nabla f\left(\boldsymbol{x}^{r}\right),\boldsymbol{x}^{r+1}-\boldsymbol{x}^{r}\right\rangle+\frac{L}{2}\left\|\boldsymbol{x}^{r+1}-\boldsymbol{x}^{r}\right\|^{2}$
$\leq f\left(\boldsymbol{x}^{r}\right)+\gamma\left\langle\nabla f\left(\boldsymbol{x}^{r}\right),{\Delta}^{r+1}_{G}\right\rangle+\frac{L\gamma^{2}}{2}\left\|{\Delta}^{r+1}_{G}\right\|^{2}$
$=f\left(\boldsymbol{x}^{r}\right)-\gamma\left\|\nabla f\left(\boldsymbol{x}^{r}\right)\odot\frac{1}{SK}\sum_{i\in S^{r}}\sum_{k=1}^{K}\vartheta_{i}^{r,k}\right\|^{2}+\gamma\left\langle\nabla f\left(\boldsymbol{x}^{r}\right),\nabla f\left(\boldsymbol{x}^{r}\right)\odot\frac{1}{SK}\sum_{i\in S^{r}}\sum_{k=1}^{K}\vartheta_{i}^{r,k}-{\Delta}^{r+1}_{G}\right\rangle$
$+\frac{L\gamma^{2}}{2}\left\|{\Delta}^{r+1}_{G}\right\|^{2}.$
Since $\boldsymbol{x}^{r+1}=\boldsymbol{x}^{r}-\gamma{\Delta}^{r+1}_{G}$, using Young’s inequality, we further have
$\mathbb{E}f\left(\boldsymbol{x}^{r+1}\right)$
$\leq f\left(\boldsymbol{x}^{r}\right)-(\gamma G_{g}^{2}-\frac{\gamma}{2})\left\|\nabla f\left(\boldsymbol{x}^{r}\right)\right\|^{2}+\frac{\gamma}{2}\left\|\nabla f\left(\boldsymbol{x}^{r}\right)\odot\frac{1}{SK}\sum_{i\in S^{r}}\sum_{k=1}^{K}\vartheta_{i}^{r,k}-{\Delta}^{r+1}_{G}\right\|^{2}$
$+L\gamma^{2}\left(\left\|\nabla f\left(\boldsymbol{x}^{r}\right)\odot\frac{1}{SK}\sum_{i\in S^{r}}\sum_{k=1}^{K}\vartheta_{i}^{r,k}\right\|^{2}+\left\|\nabla f\left(\boldsymbol{x}^{r}\right)\odot\frac{1}{SK}\sum_{i\in S^{r}}\sum_{k=1}^{K}\vartheta_{i}^{r,k}-{\Delta}^{r+1}_{G}\right\|^{2}\right)$
$\leq f\left(\boldsymbol{x}^{r}\right)-(\gamma G_{g}^{2}-\frac{\gamma}{2}-L\gamma^{2}G_{g}^{2})\left\|\nabla f\left(\boldsymbol{x}^{r}\right)\right\|^{2}+\frac{13\gamma}{24}\left\|\nabla f\left(\boldsymbol{x}^{r}\right)\odot\frac{1}{SK}\sum_{i\in S^{r}}\sum_{k=1}^{K}\vartheta_{i}^{r,k}-{\Delta}^{r+1}_{G}\right\|^{2}$
$\leq f\left(\boldsymbol{x}^{r}\right)-\frac{11\gamma}{24}\left\|\nabla f\left(\boldsymbol{x}^{r}\right)\right\|^{2}+\frac{13\gamma}{24}\left\|\nabla f\left(\boldsymbol{x}^{r}\right)\odot\frac{1}{SK}\sum_{i\in S^{r}}\sum_{k=1}^{K}\vartheta_{i}^{r,k}-{\Delta}^{r+1}_{G}\right\|^{2},$
where the last inequality is due to $\gamma L\leq\frac{1}{24}$$(\gamma G_{g}^{2}-\frac{\gamma}{2}-L\gamma^{2}G_{g}^{2})\leq-\frac{11\gamma}{24}$. Taking the global expectation, we finish the proof. ∎
Lemma 4.
If $\gamma L\leq\frac{\alpha}{6}$, the following holds for $r\geq 1$ :
$\mathcal{E}_{r}\leq\left(1-\frac{8\alpha}{9}\right)\mathcal{E}_{r-1}+\frac{4\gamma^{2}L^{2}}{\alpha}\mathbb{E}\left[\left\|\nabla f\left(\boldsymbol{x}^{r-1}\right)\right\|^{2}\right]+\frac{2\alpha^{2}\sigma_{l}^{2}}{SK\epsilon^{2}}+4\frac{\alpha}{\epsilon^{2}}L^{2}U_{r}.$
Additionally, it holds for $r=0$ that
$\mathcal{E}_{0}\leq(1-\alpha)\mathcal{E}_{-1}+\frac{2\alpha^{2}\sigma_{l}^{2}}{SK}+4\alpha L^{2}U_{0}.$
Proof.
For $r>1$,
$\mathcal{E}_{r}=$ $\mathbb{E}\left[\left\|\frac{1}{SK}\sum_{i\in S^{r}}\sum_{k=1}^{K}\nabla f\left(\boldsymbol{x}^{r}\right)\odot\vartheta_{i}^{r,k}-{\Delta}^{r+1}_{G}\right\|^{2}\right]$
$=$ $\mathbb{E}\left[\left\|(1-\alpha)\left(\frac{1}{SK}\sum_{i\in S^{r}}\sum_{k=1}^{K}\nabla f\left(\boldsymbol{x}^{r}\right)\odot\vartheta_{i}^{r,k}-{\Delta}^{r}_{G}\right)+\alpha\left(\frac{1}{SK}\sum_{i\in S^{r}}\sum_{k=1}^{K}\nabla f\left(\boldsymbol{x}^{r}\right)\odot\vartheta_{i}^{r,k}-\frac{1}{SK}\sum_{i\in S^{r}}\sum_{k=1}^{K}g_{i}^{r,k}\odot\vartheta_{i}^{r,k}\right)\right\|^{2}\right]$
$\leq$ $\mathbb{E}\left[\left\|(1-\alpha)\left(\frac{1}{SK}\sum_{i\in S^{r}}\sum_{k=1}^{K}\nabla f\left(\boldsymbol{x}^{r}\right)\odot\vartheta_{i}^{r,k}-{\Delta}^{r}_{G}\right)\right\|^{2}\right]+\alpha^{2}\frac{1}{\epsilon^{2}}\mathbb{E}\left[\left\|\nabla f\left(\boldsymbol{x}^{r}\right)-\frac{1}{SK}\sum_{i\in S^{r}}\sum_{k=1}^{K}g_{i}^{r,k}\right\|^{2}\right]$
$+2\alpha\mathbb{E}\left[\left\langle(1-\alpha)\left(\frac{1}{SK}\sum_{i\in S^{r}}\sum_{k=1}^{K}\nabla f\left(\boldsymbol{x}^{r}\right)\odot\vartheta_{i}^{r,k}-{\Delta}^{r}_{G}\right),\frac{1}{SK}\sum_{i\in S^{r}}\sum_{k=1}^{K}\nabla f\left(\boldsymbol{x}^{r}\right)\odot\vartheta_{i}^{r,k}-\frac{1}{SK}\sum_{i\in S^{r}}\sum_{k=1}^{K}g_{i}^{r,k}\odot\vartheta_{i}^{r,k}\right\rangle\right].$
Note that $\left\{\nabla F\left(\boldsymbol{x}_{i}^{r,k};\xi_{i}^{r,k}\right)\right\}_{0\leq k<K}$ are sequentially correlated. Applying the AM-GM inequality and Lemma 1, we have
$\mathcal{E}_{r}\leq\left(1+\frac{\alpha}{2}\right)\mathbb{E}\left[\left\|(1-\alpha)\left(\nabla f\left(\boldsymbol{x}^{r}\right)-{\Delta}^{r}_{G}\right)\right\|^{2}\right]+2\frac{\alpha}{\epsilon^{2}}L^{2}U_{r}+2\frac{\alpha^{2}}{\epsilon^{2}}\left(\frac{\sigma_{l}^{2}}{SK}+L^{2}U_{r}\right).$
Using the AM-GM inequality again and Assumption 1, we have
$\mathcal{E}_{r}$ $\leq(1-\alpha)^{2}\left(1+\frac{\alpha}{2}\right)\left[\left(1+\frac{\alpha}{2}\right)\mathcal{E}_{r-1}+\left(1+\frac{2}{\alpha}\right)L^{2}\mathbb{E}\left[\left\|\boldsymbol{x}^{r}-\boldsymbol{x}^{r-1}\right\|^{2}\right]\right]+\frac{2\alpha^{2}\sigma_{l}^{2}}{SK\epsilon^{2}}+4\frac{\alpha}{\epsilon^{2}}L^{2}U_{r}$
$\leq(1-\alpha)\mathcal{E}_{r-1}+\frac{2}{\alpha}L^{2}\mathbb{E}\left[\left\|\boldsymbol{x}^{r}-\boldsymbol{x}^{r-1}\right\|^{2}\right]+\frac{2\alpha^{2}\sigma_{l}^{2}}{\epsilon^{2}SK}+4\frac{\alpha}{\epsilon^{2}}L^{2}U_{r}$
$\leq\left(1-\frac{8\alpha}{9}\right)\mathcal{E}_{r-1}+4\frac{\gamma^{2}L^{2}}{\alpha}\mathbb{E}\left[\left\|\nabla f\left(\boldsymbol{x}^{r-1}\right)\right\|^{2}\right]+\frac{2\alpha^{2}\sigma_{l}^{2}}{\epsilon^{2}SK}+4\frac{\alpha}{\epsilon^{2}}L^{2}U_{r},$
where we plug in $\left\|\boldsymbol{x}^{r}-\boldsymbol{x}^{r-1}\right\|^{2}\leq 2\gamma^{2}\left(\left\|\nabla f\left(\boldsymbol{x}^{r-1}\right)\right\|^{2}+\left\|{\Delta}^{r}_{G}-\nabla f\left(\boldsymbol{x}^{r-1}\right)\right\|^{2}\right)$ and use $\gamma L\leq\frac{\alpha}{6}$ in the last inequality. Similarly for $r=0$,
$\mathcal{E}_{0}$ $\leq\left(1+\frac{\alpha}{2}\right)\mathbb{E}\left[\left\|(1-\alpha)\left(\nabla f\left(\boldsymbol{x}^{0}\right)-g^{0}\right)\right\|^{2}\right]+2\frac{\alpha}{\epsilon^{2}}L^{2}U_{0}+2\frac{\alpha^{2}}{\epsilon^{2}}\left(\frac{\sigma_{l}^{2}}{SK}+L^{2}U_{0}\right)$
$\leq(1-\alpha)\mathcal{E}_{-1}+\frac{2\alpha^{2}\sigma_{l}^{2}}{SK\epsilon^{2}}+4\frac{\alpha}{\epsilon^{2}}L^{2}U_{0}.$
Lemma 5.
If $\eta LK\leq\frac{1}{\alpha}$, the following holds for $r\geq 0$ :
$U_{r}\leq 2eK^{2}\Xi_{r}+K\eta^{2}\alpha^{2}\frac{1}{\epsilon^{2}}\sigma_{l}^{2}\left(1+2K^{3}L^{2}\eta^{2}\alpha^{2}\right).$
Proof.
Recall that $\xi_{i}^{r,k}:=\mathbb{E}\left[\boldsymbol{x}_{i}^{r,k+1}-\boldsymbol{x}_{i}^{r,k}\mid\mathcal{F}_{i}^{r,k}\right]=-\eta\left((1-\alpha){\Delta}^{r}_{G}+\alpha\nabla f_{i}\left(\boldsymbol{x}_{i}^{r,k}\right)\odot\vartheta_{i}^{r,k}\right)$. Then we have
$\mathbb{E}\left[\left\|\xi_{i}^{r,j}-\xi_{i}^{r,j-1}\right\|^{2}\right]$ $\leq\frac{1}{\epsilon^{2}}\eta^{2}L^{2}\alpha^{2}\mathbb{E}\left[\left\|x_{i}^{r,j}-x_{i}^{r,j-1}\right\|^{2}\right]$
$\leq\frac{1}{\epsilon^{2}}\eta^{2}L^{2}\alpha^{2}\left(\eta^{2}\alpha^{2}\sigma_{l}^{2}+\mathbb{E}\left[\left\|\xi_{i}^{r,j-1}\right\|^{2}\right)\right.$.
For any $1\leq j\leq k-1\leq K-2$, using $\eta L\leq\frac{1}{\alpha K}\leq\frac{1}{\alpha(k+1)}$, we have
$\mathbb{E}\left[\left\|\xi_{i}^{r,j}\right\|^{2}\right]$ $\leq\left(1+\frac{1}{k}\right)\mathbb{E}\left[\left\|\xi_{i}^{r,j-1}\right\|^{2}\right]+(1+k)\mathbb{E}\left[\left\|\xi_{i}^{r,j}-\xi_{i}^{r,j-1}\right\|^{2}\right]$
$\leq\left(1+\frac{2}{k}\right)\mathbb{E}\left[\left\|\xi_{i}^{r,j-1}\right\|^{2}\right]+(k+1)\frac{1}{\epsilon^{2}}L^{2}\eta^{4}\alpha^{4}\sigma_{l}^{2}$
$\leq e^{2}\mathbb{E}\left[\left\|\xi_{i}^{r,0}\right\|^{2}\right]+4\frac{1}{\epsilon^{2}}k^{2}L^{2}\eta^{4}\alpha^{4}\sigma_{l}^{2}.$
where the last inequality is by unrolling the recursive bound and using $\left(1+\frac{2}{k}\right)^{k}\leq e^{2}$. By Lemma 1 , it holds that for $k\geq 2$,
$\mathbb{E}\left[\left\|\boldsymbol{x}_{i}^{r,k}-\boldsymbol{x}^{r}\right\|^{2}\right]$ $\leq 2\mathbb{E}\left[\left\|\sum_{j=0}^{k-1}\xi_{i}^{r,j}\right\|^{2}\right]+2\frac{1}{\epsilon^{2}}k\eta^{2}\alpha^{2}\sigma_{l}^{2}$
$\leq 2k\sum_{j=0}^{k-1}\mathbb{E}\left[\left\|\xi_{i}^{r,k}\right\|^{2}\right]+2\frac{1}{\epsilon^{2}}k\eta^{2}\alpha^{2}\sigma_{l}^{2}$
$\leq 2e^{2}k^{2}\mathbb{E}\left[\left\|\xi_{i}^{r,0}\right\|^{2}\right]+2\frac{1}{\epsilon^{2}}k\eta^{2}\alpha^{2}\sigma_{l}^{2}\left(1+4k^{3}L^{2}\eta^{2}\alpha^{2}\right).$
This is also valid for $k=0,1$. Summing up over $i$ and $k$, we finish the proof. ∎
Lemma 6.
If $288e(\eta KL)^{2}\left((1-\alpha)^{2}+e(\alpha\gamma LR)^{2}\right)\leq 1$, then it holds for $r\geq 0$ that
$\sum_{r=0}^{R-1}\Xi_{r}\leq\frac{1}{72eK^{2}L^{2}}\sum_{r=-1}^{R-2}\left(\mathcal{E}_{r}+\mathbb{E}\left[\left\|\nabla f\left(\boldsymbol{x}^{r}\right)\right\|^{2}\right]\right)+2\eta^{2}\alpha^{2}\frac{1}{\epsilon^{2}}eRG_{0}.$
Proof.
Note that $\xi_{i}^{r,0}=-\eta\left((1-\alpha){\Delta}^{r}_{G}+\alpha\nabla f_{i}\left(\boldsymbol{x}^{r}\right)\odot\vartheta_{i}^{r,k}\right)$,
$\frac{1}{N}\sum_{i=1}^{N}\left\|\xi_{i}^{r,0}\right\|^{2}\leq 2\eta^{2}\left((1-\alpha)^{2}\left\|{\Delta}^{r}_{G}\right\|^{2}+\alpha^{2}\frac{1}{N}\sum_{i=1}^{N}\left\|\nabla f_{i}\left(\boldsymbol{x}^{r}\right)\right\|^{2}\right).$
Using Young’s inequality, we have for any $q>0$ that
$\mathbb{E}\left[\left\|\nabla f_{i}\left(\boldsymbol{x}^{r}\right)\right\|^{2}\right]$ $\leq(1+q)\mathbb{E}\left[\left\|\nabla f_{i}\left(\boldsymbol{x}^{r-1}\right)\right\|^{2}\right]+\left(1+q^{-1}\right)L^{2}\mathbb{E}\left[\left\|\boldsymbol{x}^{r}-\boldsymbol{x}^{r-1}\right\|^{2}\right]$
$\leq(1+q)\mathbb{E}\left[\left\|\nabla f_{i}\left(\boldsymbol{x}^{r-1}\right)\right\|^{2}\right]+2\left(1+q^{-1}\right)\gamma^{2}L^{2}\left(\mathcal{E}_{r-1}+\mathbb{E}\left[\left\|\nabla f\left(\boldsymbol{x}^{r-1}\right)\right\|^{2}\right]\right)$
$\leq(1+q)^{r}\mathbb{E}\left[\left\|\nabla f_{i}\left(\boldsymbol{x}^{0}\right)\right\|^{2}\right]+\frac{2}{q}\gamma^{2}L^{2}\sum_{j=0}^{r-1}\left(\mathcal{E}_{j}+\mathbb{E}\left[\left\|\nabla f\left(x^{j}\right)\right\|^{2}\right)(1+q)^{r-j}\right.$.
Take $q=\frac{1}{r}$ and we have
$\mathbb{E}\left[\left\|\nabla f_{i}\left(\boldsymbol{x}^{r}\right)\right\|^{2}\right]\leq e\mathbb{E}\left[\left\|\nabla f_{i}\left(\boldsymbol{x}^{0}\right)\right\|^{2}\right]+2e(r+1)\gamma^{2}L^{2}\sum_{j=0}^{r-1}\left(\mathcal{E}_{j}+\mathbb{E}\left[\left\|\nabla f\left(x^{j}\right)\right\|^{2}\right)\right.$. (8)
Note that this inequality is valid for $r=0$. Therefore, using (8), we have
$\sum_{r=0}^{R-1}\Xi_{r}\leq$ $\sum_{r=0}^{R-1}2\eta^{2}\mathbb{E}\left[(1-\alpha)^{2}\left\|{\Delta}^{r}_{G}\right\|^{2}+\alpha^{2}\frac{1}{\epsilon^{2}}\frac{1}{N}\sum_{i=1}^{N}\left\|\nabla f_{i}\left(\boldsymbol{x}^{r}\right)\right\|^{2}\right]$
$\leq$ $\sum_{r=0}^{R-1}2\eta^{2}\left(2(1-\alpha)^{2}\left(\mathcal{E}_{r-1}+\mathbb{E}\left[\left\|\nabla f\left(\boldsymbol{x}^{r-1}\right)\right\|^{2}\right]\right)+\alpha^{2}\frac{1}{\epsilon^{2}}\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\left\|\nabla f_{i}\left(\boldsymbol{x}^{r}\right)\right\|^{2}\right]\right)$
$\leq$ $\sum_{r=0}^{R-1}4\eta^{2}(1-\alpha)^{2}\left(\mathcal{E}_{r-1}+\mathbb{E}\left[\left\|\nabla f\left(\boldsymbol{x}^{r-1}\right)\right\|^{2}\right]\right)$
$+2\eta^{2}\alpha^{2}\frac{1}{\epsilon^{2}}\sum_{r=0}^{R-1}\left(\frac{e}{N}\sum_{i=1}^{N}\mathbb{E}\left[\left\|\nabla f_{i}\left(\boldsymbol{x}^{0}\right)\right\|^{2}\right]+2e(r+1)(\gamma L)^{2}\sum_{j=0}^{r-1}\left(\mathcal{E}_{j}+\mathbb{E}\left[\left\|\nabla f\left(x^{j}\right)\right\|^{2}\right]\right)\right)$
$\leq$ $4\eta^{2}(1-\alpha)^{2}\sum_{r=0}^{R-1}\left(\mathcal{E}_{r-1}+\mathbb{E}\left[\left\|\nabla f\left(\boldsymbol{x}^{r-1}\right)\right\|^{2}\right]\right)$
$+2\eta^{2}\alpha^{2}\frac{1}{\epsilon^{2}}\left(eRG_{0}+2e(\gamma LR)^{2}\sum_{r=0}^{R-2}\left(\mathcal{E}_{r}+\mathbb{E}\left[\left\|\nabla f\left(\boldsymbol{x}^{r}\right)\right\|^{2}\right]\right)\right).$
Rearranging the equation and applying the upper bound of $\eta$, we finish the proof. ∎
Theorem 3 (Convergence for non-convex functions).
Under Assumption 1 and 2 and 3 , if we take $g^{0}=0$,
$\alpha=\min\left\{,\sqrt{\frac{SKL\Delta\epsilon^{2}}{\sigma_{l}^{2}R}}\right\}\text{ for any constant }c\in(0,1],\quad\gamma=\min\left\{\frac{1}{24LG_{g}},\frac{\alpha}{6L}\right\},$
$\eta KL\lesssim\min\left\{1,\frac{1}{\alpha\gamma LR},\left(\frac{L\Delta}{G_{0}\alpha^{3}R}\right)^{1/2},\frac{1}{(\alpha N)^{1/2}},\frac{1}{\left(\alpha^{3}NK\right)^{1/4}}\right\}.$
then FedAdamW converges as
$\frac{1}{R}\sum_{r=0}^{R-1}\mathbb{E}\left[\left\|\nabla f\left(\boldsymbol{x}^{r}\right)\right\|^{2}\right]\lesssim\sqrt{\frac{L\Delta\sigma_{l}^{2}}{SKR}}+\frac{L\Delta}{R}.$
Here $G_{0}:=\frac{1}{N}\sum_{i=1}^{N}\left\|\nabla f_{i}\left(\boldsymbol{x}^{0}\right)\right\|^{2}$.
Proof.
Combining Lemma 3 and 4, we have
$\mathcal{E}_{r}\leq$ $\left(1-\frac{8\alpha}{9}\right)\mathcal{E}_{r-1}+4\frac{(\gamma L)^{2}}{\alpha}\mathbb{E}\left[\left\|\nabla f\left(\boldsymbol{x}^{r-1}\right)\right\|^{2}\right]+\frac{2\alpha^{2}\sigma_{l}^{2}}{SK\epsilon^{2}}$
$+4\frac{\alpha}{\epsilon^{2}}L^{2}\left(2eK^{2}\Xi_{r}+K\eta^{2}\alpha^{2}\frac{1}{\epsilon^{2}}\sigma_{l}^{2}\left(1+2K^{3}L^{2}\eta^{2}\alpha^{2}\right)\right$.,
and
$\mathcal{E}_{0}\leq(1-\alpha)\mathcal{E}_{-1}+\frac{2\alpha^{2}\sigma_{l}^{2}}{SK\epsilon^{2}}+4\alpha\frac{1}{\epsilon^{2}}L^{2}\left(2eK^{2}\Xi_{0}+K\eta^{2}\alpha^{2}\frac{1}{\epsilon^{2}}\sigma_{l}^{2}\left(1+2K^{3}L^{2}\eta^{2}\alpha^{2}\right)\right).$
Summing over $r$ from 0 to R1R-1 and applying Lemma 6,
$\sum_{r=0}^{R-1}\mathcal{E}_{r}\leq$ $\left(1-\frac{8\alpha}{9}\right)\sum_{r=-1}^{R-2}\mathcal{E}_{r}+4\frac{(\gamma L)^{2}}{\alpha}\sum_{r=0}^{R-2}\mathbb{E}\left[\left\|\nabla f\left(\boldsymbol{x}^{r}\right)\right\|^{2}\right]+2\frac{\alpha^{2}\sigma_{l}^{2}}{SK\epsilon^{2}}R$
$+4\alpha\frac{1}{\epsilon^{2}}L^{2}\left(2eK^{2}\sum_{r=0}^{R-1}\Xi_{r}+RK\eta^{2}\alpha^{2}\frac{1}{\epsilon^{2}}\sigma_{l}^{2}\left(1+2K^{3}L^{2}\eta^{2}\alpha^{2}\right)\right)$
[…]
$+\frac{2\alpha^{2}\sigma_{l}^{2}}{SK\epsilon^{2}}R+4\alpha^{3}\frac{1}{\epsilon^{4}}(\eta KL)^{2}\left(\frac{1}{K}+2(\eta KL\alpha)^{2}\right)\sigma_{l}^{2}R$
[…]
Here in the last inequality we apply
$4\alpha\frac{1}{\epsilon^{4}}(\eta KL)^{2}\left(\frac{1}{K}+2(\eta KL\alpha)^{2}\right)\leq\frac{2}{NK}\quad\text{ and }\quad\gamma L\leq\frac{\alpha}{6}.$
Therefore,
$\sum_{r=0}^{R-1}\mathcal{E}_{r}\leq\frac{9}{7\alpha}\mathcal{E}_{-1}+\frac{2}{7}\mathbb{E}\left[\sum_{r=-1}^{R-2}\left\|\nabla f\left(\boldsymbol{x}^{r}\right)\right\|^{2}\right]+\frac{144}{7}\frac{1}{\epsilon^{4}}(e\alpha\eta KL)^{2}G_{0}R+\frac{36\alpha\sigma_{l}^{2}}{7SK\epsilon^{2}}R.$
Combine this inequality with Lemma 3 and we get
$\frac{1}{\gamma}\mathbb{E}\left[f\left(\boldsymbol{x}^{r}\right)-f\left(\boldsymbol{x}^{0}\right)\right]\leq-\frac{1}{7}\sum_{r=0}^{R-1}\mathbb{E}\left[\left\|\nabla f\left(\boldsymbol{x}^{r}\right)\right\|^{2}\right]+\frac{39}{56\alpha}\mathcal{E}_{-1}+\frac{78}{7}\frac{1}{\epsilon^{4}}(e\alpha\eta KL)^{2}G_{0}R+\frac{39\alpha\sigma_{l}^{2}}{14SK\epsilon^{2}}R.$
Finally, noticing that $g^{0}=0$ implies $\mathcal{E}_{-1}\leq 2L\left(f\left(\boldsymbol{x}^{0}\right)-f^{*}\right)=2L\Delta$, we obtain
$\frac{1}{R}\sum_{r=0}^{R-1}\mathbb{E}\left[\left\|\nabla f\left(\boldsymbol{x}^{r}\right)\right\|^{2}\right]$ $\lesssim\frac{L\Delta}{\gamma LR}+\frac{\mathcal{E}_{-1}}{\alpha R}+(\alpha\eta KL)^{2}\frac{1}{\epsilon^{4}}G_{0}+\frac{\alpha\sigma_{l}^{2}}{SK\epsilon^{2}}$
$\lesssim\frac{L\Delta}{R}+\frac{L\Delta}{\alpha R}+\frac{\alpha\sigma_{l}^{2}}{SK\epsilon^{2}}+(\alpha\eta KL)^{2}G_{0}\frac{1}{\epsilon^{4}}$
$\lesssim\frac{L\Delta}{R}+\sqrt{\frac{L\Delta\sigma_{l}^{2}}{SKR\epsilon^{2}}}.$
Appendix B Appendix B: Generalization Analysis
Analysis on hypothesis posterior. According to the classical PAC-Bayesian theory (McAllester 1998), the generalization error bound is closely related to the hypothesis posterior learned by an algorithm. We therefore begin by analyzing the hypothesis posterior induced by FedAdamW, followed by its impact on generalization. Following prior work (Xie et al. 2021; Mandt, Hoffman, and Blei 2016; Chaudhari and Soatto 2018; Jastrzbski et al. 2017; Zhu et al. 2018; Zhou et al. 2020, 2024), we adopt the stochastic differential equation (SDE) view of optimization, as it captures similar convergence behavior to that of the actual algorithm. We focus on a simplified version of the second-moment estimate used in AdaGrad (Duchi, Hazan, and Singer 2011), which is given by: $\mathrm{diag}\left(\boldsymbol{n}_{r+1}\right)$, which is commonly used in practice. The observed fisher information matrix, $\frac{1}{n}\sum_{i=1}^{n}\nabla F\left(\boldsymbol{x}^{r};\boldsymbol{\xi}_{i}\right)\nabla F\left(\boldsymbol{x}^{r};\boldsymbol{\xi}_{i}\right)^{\top}$ serves as a reliable approximation of the local Hessian matrix $\boldsymbol{H}_{\boldsymbol{x}^{r}}$ (Pawitan 2001; Jastrzbski et al. 2017; Zhu et al. 2018). Based on these insights, we follow previous works and treat $\boldsymbol{n}_{r}$ as a proxy for the Hessian matrix in our analysis.
$\boldsymbol{n}_{r}\approx\boldsymbol{H}_{\boldsymbol{x}^{r}}.$
Accordingly, the updating rule of FedAdamW can be formulated as follows:
$\boldsymbol{x}^{r+1}$ $=\boldsymbol{x}^{r}-\eta\frac{1}{SK\eta}\sum_{i=1}^{S}\Delta\boldsymbol{x}_{i}^{r}$ (9)
$=\!\boldsymbol{x}^{r}\!-\!\eta\frac{1}{SK\eta}\sum_{i=1}^{S}\sum_{i=1}^{K}\!-\eta\Big(\hat{\boldsymbol{m}}^{r,k}_{i}\odot\vartheta_{i}^{r,k}\!-\!w\boldsymbol{x}_{i}^{r,k}\!+\!\alpha\boldsymbol{\Delta}_{G}\Big)$
$\approx\boldsymbol{x}^{r}-\eta\boldsymbol{Q}_{r}\nabla F\left(\boldsymbol{x}^{r}\right)-\eta\lambda\boldsymbol{x}^{r}+\eta\boldsymbol{Q}_{r}\boldsymbol{u}^{r}.$
Let $\boldsymbol{u}^{r}=\nabla F(\boldsymbol{x}^{r})-\frac{1}{N}\sum_{i=1}^{N}\nabla F_{i}(\boldsymbol{x}^{r})$ denote the heterogeneity gradient noise, and define $\boldsymbol{Q}_{r}=\boldsymbol{n}^{-1/2}$ for simplicity. In the FedAdamW formulation, we omit the small constant $\epsilon$ for clarity, as it does not affect subsequent results. Following (Mandt, Hoffman, and Blei 2017; Smith and Le 2017; Chaudhari and Soatto 2018; Zhu et al. 2018), we assume the noise $\boldsymbol{u}^{r}$ follows a Gaussian distribution $\mathcal{N}(\mathbf{0},\boldsymbol{C}_{\boldsymbol{x}^{r}})$ by the Central Limit Theorem. Under this assumption, the SDE of FedAdamW can be expressed as:
$\mathrm{d}\boldsymbol{x}^{r}=-\boldsymbol{Q}_{r}\nabla F\left(\boldsymbol{x}^{r}\right)\mathrm{d}r-\lambda\boldsymbol{x}^{r}\mathrm{~d}r+\boldsymbol{Q}_{r}\left(2\boldsymbol{\Sigma}_{r}\right)^{\frac{1}{2}}\mathrm{~d}\xi_{r},$
where $\mathrm{d}\xi_{r}\sim\mathcal{N}(0,\boldsymbol{I}\mathrm{~d}r)$ and $\boldsymbol{\Sigma}_{r}=\frac{\eta}{2}\boldsymbol{C}_{\boldsymbol{x}^{r}}$. Here, $\boldsymbol{C}_{\boldsymbol{x}^{r}}$ is defined in the above literatures as
$\boldsymbol{C}_{\boldsymbol{x}^{r}}\!\!=\!\!\frac{1}{S}\!\left[\!\frac{1}{N}\sum_{i=1}^{n}\nabla f\left(\boldsymbol{x}^{r};\boldsymbol{\xi}_{i}\right)\!\nabla f\left(\boldsymbol{x}^{r};\boldsymbol{\xi}_{i}\right)^{\top}\!\!-\!\!\nabla F\left(\boldsymbol{x}^{r}\right)\!\nabla F\left(\boldsymbol{x}^{r}\right)^{\top}\!\!\right],$
where $N$ is the number of all clients, and $S$ is participate clients. Since we analyze the local convergence around an optimum, this means $\nabla F\left(\boldsymbol{x}^{r}\right)\approx 0$ and the variance of the gradient noise would dominate (Mandt, Hoffman, and Blei 2017; Smith and Le 2017; Chaudhari and Soatto 2018) . So following these works, we approximate $\boldsymbol{C}_{\boldsymbol{x}^{r}}$ as
$\boldsymbol{C}_{\boldsymbol{x}^{r}}\approx\frac{1}{S}\left[\frac{1}{N}\sum_{i=1}^{N}\nabla f\left(\boldsymbol{x}^{r};\boldsymbol{\xi}_{i}\right)\nabla f\left(\boldsymbol{x}^{r};\boldsymbol{\xi}_{i}\right)^{\top}\right]\approx\frac{1}{S}\boldsymbol{H}_{\boldsymbol{x}^{r}}.$
Based on these approximations, the hypothesis posterior distribution induced by FedAdamW is characterized in Lemma 7.
Lemma 7.
Assume the objective loss can be approximated by a second-order Taylor approximation, namely, $F(\boldsymbol{x})=F\left(\boldsymbol{x}_{*}\right)+\frac{1}{2}\left(\boldsymbol{x}-\boldsymbol{x}_{*}\right)^{\top}\boldsymbol{H}_{*}\left(\boldsymbol{x}-\boldsymbol{x}_{*}\right)$. Then the solution $\boldsymbol{x}^{r}$ of FedAdamW obeys a Gaussian distribution $\mathcal{N}\left(\boldsymbol{x}_{*},\boldsymbol{M}_{\text{FedAdamW }}\right)$ where the covariance matrix $\boldsymbol{M}_{\text{FedAdamW }}=\mathbb{E}\left[\boldsymbol{x}^{r}\boldsymbol{x}^{\top}\right]$ is defined
$\boldsymbol{M}_{AdamW}=\frac{\eta}{2B}\boldsymbol{U}\left(\boldsymbol{S}^{\frac{1}{2}}+\lambda\boldsymbol{I}\right)^{-1}\boldsymbol{U}^{\top},$
where $\boldsymbol{U}\boldsymbol{S}\boldsymbol{U}^{\top}$ is the SVD of $\boldsymbol{H}_{*}$ in which $\boldsymbol{S}=\operatorname{diag}\left(\sigma_{1},\sigma_{2},\cdots,\sigma_{d}\right)$ with singular values $\left\{\sigma_{i}\right\}_{i=1}^{d}$.
Lemma  7 shows that FedAdamW converges to a solution concentrated around the minimum $\boldsymbol{x}_{*}$, confirming its favorable convergence from the SDE perspective. We further examine the effect of the decoupled weight decay parameter $\lambda$. As $\lambda$ increases, the singular values of the covariance matrix $\boldsymbol{M}_{\texttt{FedAdamW}}$ decrease, indicating improved stability. This suggests that decoupled weight decay not only enhances convergence to $\boldsymbol{x}_{*}$ but also leads to better generalization performance.
Building on the posterior analysis, we adopt the PAC-Bayesian framework (McAllester 1998) to study the generalization performance of FedAdamW. Given an algorithm $\mathcal{A}$ and a training set $\mathcal{D}_{\text{tr}}$ with samples $\boldsymbol{\xi}\sim\mathcal{D}$, the learned posterior hypothesis follows a distribution $\mathcal{P}\sim\mathcal{N}(\boldsymbol{x}_{*},\boldsymbol{M}_{\text{{FedAdamW}}})$ as shown in Lemma 7. We define the expected risk as $\mathbb{E}_{\boldsymbol{\xi}\sim\mathcal{D},\boldsymbol{x}\sim\mathcal{P}}[f(\boldsymbol{x},\boldsymbol{\xi})]$ and the empirical risk as $\mathbb{E}_{\boldsymbol{\xi}\in\mathcal{D}_{\mathrm{tr}},\boldsymbol{x}\sim\mathcal{P}}[f(\boldsymbol{x},\boldsymbol{\xi})]$. In practice, the prior is often chosen as a Gaussian $\mathcal{P}_{\text{pre}}\sim\mathcal{N}(\mathbf{0},\rho\boldsymbol{I})$ (Szegedy et al. 2015; He et al. 2016; Dosovitskiy et al. 2020), due to the lack of prior knowledge about the posterior. Based on Lemma 7, we establish the generalization bound in Theorem 4.
Theorem 4.
Assume the prior hypothesis $\boldsymbol{x}_{0}$ satisfies $\mathcal{P}_{\text{pre }}\sim\mathcal{N}(\mathbf{0},\rho\boldsymbol{I})$. Then the expected risk for the posterior hypothesis $\boldsymbol{x}\sim\mathcal{P}$ of FedAdamW learned on training dataset $\mathcal{D}_{tr}\sim\mathcal{D}$ with $n$ samples holds
$\mathbb{E}_{\boldsymbol{\xi}\sim\mathcal{D},\boldsymbol{x}\sim\mathcal{P}}[f(\boldsymbol{x},\boldsymbol{\xi})]-\mathbb{E}_{\boldsymbol{\xi}\in\mathcal{D}_{tr},\boldsymbol{x}\sim\mathcal{P}}[f(\boldsymbol{x},\boldsymbol{\xi})]\leq$ (10)
$\frac{\sqrt{8}}{\sqrt{n}}\left(\sum_{i=1}^{d}\log\frac{2\rho b\left(\sigma_{i}^{\frac{1}{2}}+\lambda\right)}{\eta}\!+\!\frac{\eta}{2\rho b}\sum_{i=1}^{d}\frac{1}{\sigma_{i}^{\frac{1}{2}}+\lambda}\!+\!c_{0}\!\right)^{\frac{1}{2}},$
with at least probability $1-\tau$, where $\tau\in(0,1)$ and $c_{0}=\frac{1}{2\rho}\left\|\boldsymbol{x}_{*}\right\|^{2}-\frac{d}{2}+2\ln\left(\frac{2n}{\tau}\right)$.
Theorem 2 shows that the generalization error of FedAdamW can be upper bounded by $\mathcal{O}(1/\sqrt{n})$ (up to constant factors), consistent with classical results from PAC theory, stability, and uniform convergence (Vapnik 2006). We further analyze the impact of the decoupled weight decay parameter $\lambda$ on this bound. As $\lambda$ increases, the first term $\sum_{i=1}^{d}\log 2\rho b\left(\sigma_{i}^{1/2}+\lambda\right)\eta^{-1}$ increases, while the second term $\frac{\eta}{2\rho b}\sum_{i=1}^{d}\left(\sigma_{i}^{1/2}+\lambda\right)^{-1}$ decreases. Although choosing the optimal $\lambda$ is challenging in practice, this trade-off suggests that tuning $\lambda$ appropriately can lead to a smaller generalization error. Since the empirical risk $\mathbb{E}_{\boldsymbol{\xi}\in\mathcal{D}_{\mathrm{tr}},\boldsymbol{x}\sim\mathcal{P}}[f(\boldsymbol{x},\boldsymbol{\xi})]$ is typically small for modern over-parameterized models, adjusting $\lambda$ helps reduce the expected risk. This explains why FedAdamW often outperforms Local Adam (which corresponds to $\lambda=0$). To the best of our knowledge, this is the first theoretical result that explicitly demonstrates how weight decay improves the generalization in FL.
To begin with, we first give one useful lemma to prove our generalization error bound.
Lemma 8 (PAC-Bayesian generalization bound).
(McAllester, 1999) For any $\tau\in(0,1)$, the expected risk for the posterior hypothesis of an algorithm over a training dataset $\mathcal{D}_{tr}\sim\mathcal{D}$ with $n$ samples holds with at least probability $1-\tau$ :
$\mathbb{E}_{\boldsymbol{\xi}\sim\mathcal{D},\boldsymbol{x}\sim\mathcal{P}}[f(\boldsymbol{x},\boldsymbol{\xi})]-\mathbb{E}_{\boldsymbol{\xi}\in\mathcal{D}_{tr},\boldsymbol{x}\sim\mathcal{P}}[f(\boldsymbol{x},\boldsymbol{\xi})]\leq 4\sqrt{\frac{1}{n}\left(KL\left(\mathcal{P}\|\mathcal{P}_{pre}\right)+\ln\left(\frac{2n}{\tau}\right)\right)},$
[…]
Proof of Lemma 7
Proof.
Based on the assumptions in Lemma 7 and the inequality (9), we can write the SDE equations as follows:
$\mathrm{d}x^{r}$ $=-\boldsymbol{Q}_{r}\nabla F\left(\boldsymbol{x}^{r}\right)\mathrm{d}r-\lambda\boldsymbol{x}^{r}\mathrm{~d}r+\boldsymbol{Q}_{r}\left(2\boldsymbol{\Sigma}_{r}\right)^{\frac{1}{2}}\mathrm{~d}\zeta_{r}$
$=-\boldsymbol{Q}_{r}\boldsymbol{H}_{*}\boldsymbol{x}^{r}\mathrm{~d}r-\lambda\boldsymbol{x}^{r}\mathrm{~d}r+\boldsymbol{Q}_{r}\left(2\boldsymbol{\Sigma}_{r}\right)^{\frac{1}{2}}\mathrm{~d}\zeta_{r}$
$=-\left(\boldsymbol{H}_{*}^{\frac{1}{2}}+\lambda\boldsymbol{I}\right)\boldsymbol{x}^{r}\mathrm{~d}r+\sqrt{\frac{\eta}{b}}\mathrm{~d}\zeta_{r},$
where $\mathrm{d}\zeta_{r}\sim\mathcal{N}(0,\boldsymbol{I}\mathrm{~d}r),\boldsymbol{\Sigma}_{r}\approx\frac{\eta}{2B}\boldsymbol{H}_{*};\boldsymbol{Q}_{r}=\boldsymbol{H}_{*}^{-\frac{1}{2}}$. Then for this Ornstein-Uhlenbeck process, we can compute its closed form solution as follows:
$\boldsymbol{x}^{r}=\exp\left(-\left(\boldsymbol{H}_{*}^{\frac{1}{2}}+\lambda\boldsymbol{I}\right)r\right)\boldsymbol{x}_{0}+\sqrt{\frac{\eta}{b}}\int_{0}^{t}\exp\left(-\left(\boldsymbol{H}_{*}^{\frac{1}{2}}+\lambda\boldsymbol{I}\right)\left(r-r^{\prime}\right)\right)\mathrm{d}\zeta_{r^{\prime}}.$
Let $\boldsymbol{M}=\mathbb{E}\left[\boldsymbol{x}^{r}\boldsymbol{x}^{r}\right]$. In this way, considering $\left(\boldsymbol{H}_{*}^{\frac{1}{2}}+\lambda\boldsymbol{I}\right)=\left(\boldsymbol{H}_{*}^{\frac{1}{2}}+\lambda\boldsymbol{I}\right)^{\top}$, we follow (Mandt et al., 2017) (see their Appendix b) and can further compute the algebraic relation for the stationary covariance of the multivariate Ornstein-Uhlenbeck process as follows:
$\left(\boldsymbol{H}_{*}^{\frac{1}{2}}+\lambda\boldsymbol{I}\right)\boldsymbol{M}+\boldsymbol{M}\left(\boldsymbol{H}_{*}^{\frac{1}{2}}+\lambda\boldsymbol{I}\right)$
$=$ $\frac{\eta}{b}\int_{-\infty}^{r}\left(\boldsymbol{H}_{*}^{\frac{1}{2}}+\lambda\boldsymbol{I}\right)\exp\left(-\left(\boldsymbol{H}_{*}^{\frac{1}{2}}+\lambda\boldsymbol{I}\right)\left(r-r^{\prime}\right)\right)\exp\left(-\left(\boldsymbol{H}_{*}^{\frac{1}{2}}+\lambda\boldsymbol{I}\right)\left(r-r^{\prime}\right)\right)\mathrm{d}r^{\prime}$
$+\frac{\eta}{b}\int_{-\infty}^{r}\exp\left(-\left(\boldsymbol{H}_{*}^{\frac{1}{2}}+\lambda\boldsymbol{I}\right)\left(r-r^{\prime}\right)\right)\exp\left(-\left(\boldsymbol{H}_{*}^{\frac{1}{2}}+\lambda\boldsymbol{I}\right)\left(r-r^{\prime}\right)\right)\mathrm{d}r^{\prime}\left(\boldsymbol{H}_{*}^{\frac{1}{2}}+\lambda\boldsymbol{I}\right)$
$=$ $\frac{\eta}{b}\int_{-\infty}^{r}\frac{\mathrm{~d}}{\mathrm{~d}r^{\prime}}\left(\exp\left(-\left(\boldsymbol{H}_{*}^{\frac{1}{2}}+\lambda\boldsymbol{I}\right)\left(r-r^{\prime}\right)\right)\exp\left(-\left(\boldsymbol{H}_{*}^{\frac{1}{2}}+\lambda\boldsymbol{I}\right)\left(r-r^{\prime}\right)\right)\right)$
$=$ $\frac{\eta}{b}\boldsymbol{I},$
where we use the lower limits of the integral vanishes by the positivity of the eigenvalues of $\boldsymbol{H}_{*}^{\frac{1}{2}}+$ $\lambda\boldsymbol{I}$. Next, let $\boldsymbol{U}\boldsymbol{S}\boldsymbol{U}^{\top}$ is the SVD of $\boldsymbol{H}_{*}$, where $\boldsymbol{S}=\operatorname{diag}\left(\sigma_{1},\sigma_{2},\cdots,\sigma_{d}\right)$. Then we have
$\boldsymbol{U}\left(\boldsymbol{S}^{\frac{1}{2}}+\lambda\boldsymbol{I}\right)\boldsymbol{U}^{\top}\boldsymbol{M}+\boldsymbol{M}\boldsymbol{U}\left(\boldsymbol{S}^{\frac{1}{2}}+\lambda\boldsymbol{I}\right)\boldsymbol{U}^{\top}=\frac{\eta}{b}\boldsymbol{I}.$
Then multiplying $\boldsymbol{U}^{\top}$ from left side and also $\boldsymbol{U}$ from right side gives
$\left(\boldsymbol{S}^{\frac{1}{2}}+\lambda\boldsymbol{I}\right)\boldsymbol{U}^{\top}\boldsymbol{M}\boldsymbol{U}+\boldsymbol{U}^{\top}\boldsymbol{M}\boldsymbol{U}\left(\boldsymbol{S}^{\frac{1}{2}}+\lambda\boldsymbol{I}\right)=\frac{\eta}{b}\boldsymbol{I}.$
Therefore, we know
$\boldsymbol{M}_{\mathrm{FedAdamW}}=\frac{\eta}{2B}\boldsymbol{U}\left(\boldsymbol{S}^{\frac{1}{2}}+\lambda\boldsymbol{I}\right)^{-1}\boldsymbol{U}^{\top}.$
We complete the proof. ∎
Proof of Theorem 2
Proof.
According to the assumption in Theorem 5 and Lemma 4, we know that for AdamW, its prior and posterior distributions are both Gaussian distribution, namely $\mathcal{P}_{\text{pre }}\sim\mathcal{N}(0,\rho\boldsymbol{I})$ and $\mathcal{P}\sim$ $\mathcal{N}\left(\boldsymbol{x}_{*},\boldsymbol{M}_{\text{FedAdamW }}\right)$ where
$\boldsymbol{M}_{\mathrm{FedAdamW}}=\frac{\eta}{2B}\boldsymbol{U}\left(\boldsymbol{S}^{\frac{1}{2}}+\lambda\boldsymbol{I}\right)^{-1}\boldsymbol{U}^{\top}.$
Here, $\boldsymbol{U}\boldsymbol{S}\boldsymbol{U}^{\top}$ is the SVD of $\boldsymbol{H}_{*}$ in which $\boldsymbol{S}=\operatorname{diag}\left(\sigma_{1},\sigma_{2},\cdots,\sigma_{d}\right)$. On the other hand, for KL between two Gaussian distributions $\boldsymbol{W}_{1}\sim\left(\boldsymbol{u}_{1},\boldsymbol{\Sigma}_{1}\right)$ and $\boldsymbol{W}_{2}\sim\left(\boldsymbol{u}_{2},\boldsymbol{\Sigma}_{2}\right)$, we can follow (Pardo, 2018) and compute it as follows:
$\operatorname{KL}\left(\boldsymbol{W}_{2}\|\boldsymbol{W}_{1}\right)=\frac{1}{2}\left[\log\frac{\operatorname{det}\left(\boldsymbol{\Sigma}_{1}\right)}{\operatorname{det}\left(\boldsymbol{\Sigma}_{2}\right)}+\operatorname{Tr}\left(\boldsymbol{\Sigma}_{1}^{-1}\boldsymbol{\Sigma}_{2}\right)\right]+\frac{1}{2}\left(\boldsymbol{u}_{1}-\boldsymbol{u}_{2}\right)^{\top}\boldsymbol{\Sigma}_{1}^{-1}\left(\boldsymbol{u}_{1}-\boldsymbol{u}_{2}\right)-\frac{d}{2}.$
Accordingly, for FedAdamW, we can compute
$\mathrm{KL}\left(\mathcal{P}\|\mathcal{P}_{\text{pre }}\right)$ $=\frac{1}{2}\left[\log\frac{\rho^{d}}{\left(\frac{\eta}{2b}\right)^{d}\prod_{i=1}^{d}\frac{1}{\sigma_{i}^{\frac{1}{2}}+\lambda}}+\frac{\eta}{2\rho b}\sum_{i=1}^{d}\frac{1}{\sigma_{i}^{\frac{1}{2}}+\lambda}+\frac{1}{2\rho}\left\|\boldsymbol{x}_{*}\right\|^{2}-\frac{d}{2}\right]$
$=\frac{1}{2}\left[\sum_{i=1}^{d}\log\frac{2\rho b\left(\sigma_{i}^{\frac{1}{2}}+\lambda\right)}{\eta}+\frac{\eta}{2\rho b}\sum_{i=1}^{d}\frac{1}{\sigma_{i}^{\frac{1}{2}}+\lambda}+\frac{1}{2\rho}\left\|\boldsymbol{x}_{*}\right\|^{2}-\frac{d}{2}\right].$
Then by using Lemma 8, it further yields the generalization bound of FedAdamW as follows:
$\mathbb{E}_{\boldsymbol{\xi}\sim\mathcal{D},\boldsymbol{x}\sim\mathcal{P}}[f(\boldsymbol{x},\boldsymbol{\xi})]-\mathbb{E}_{\boldsymbol{\xi}\in\mathcal{D}_{\mathrm{rr}},\boldsymbol{x}\sim\mathcal{P}}[f(\boldsymbol{x},\boldsymbol{\xi})]$
$\leq$ $4\sqrt{\frac{1}{2n}\left(\sum_{i=1}^{d}\log\frac{2\rho b\left(\sigma_{i}^{\frac{1}{2}}+\lambda\right)}{\eta}+\frac{\eta}{2\rho b}\sum_{i=1}^{d}\frac{1}{\sigma_{i}^{\frac{1}{2}}+\lambda}+c_{0}\right)},$
where $c_{0}=\frac{1}{2\rho}\left\|\boldsymbol{x}_{*}\right\|^{2}-\frac{d}{2}+2\ln\left(\frac{2n}{\tau}\right)$. We complete the proof. ∎
Appendix C Appendix C: Experimental Setup
Setting for ResNet-18
Table 8: A summary of CIFAR10/100, Tiny ImageNet datasets, including number of total images, number of classes, and the size of the images in the datasets.
Datasets Total Images Class Image Size
CIFAR100 60,000 100 3 × 32 × 32
Tiny ImageNet 100,000 200 3 × 64 × 64
We evaluate our methods on two widely-used benchmark datasets in federated learning: CIFAR-100 and Tiny ImageNet.
  • CIFAR-100 (Krizhevsky, Hinton et al. 2009): Contains 100 classes with 600 color images per class at a resolution of $32\times 32$. It is a standard benchmark for evaluating federated image classification methods.
  • Tiny ImageNet: A subset of ImageNet with 200 classes and 500 images per class, providing a more challenging and high-resolution classification task.
Federated Learning Configuration
We simulate a cross-device federated learning environment using the following settings:
  • Number of clients: 100
  • Client participation rate: 10% per round
  • Communication rounds: 300
  • Local update steps: 50 iterations
  • Batch size: 50
We perform grid search to tune the learning rates for each algorithm:
  • FedAvg, SCAFFOLD, FedCM, and FedAdam use a local learning rate of 0.1.
  • Local Adam, FedLADA, Local AdamW, and FedAdamW use a local learning rate of 3e-4.
  • FedAdam uses a global learning rate of 0.01; FedAdamW uses 1.0.
Model Architecture
We adopt ResNet-18 as the backbone model. To better adapt it to CIFAR-100, we modify its architecture following standard practices (He et al. 2016):
  • Replace the original $7\times 7$ convolution with a $3\times 3$ kernel.
  • Remove the initial downsampling layers (stride-2 convolution and max pooling).
We also compare Batch Normalization (BN) and Group Normalization (GN) in ResNet-18. Empirically, BN outperforms GN on CIFAR-100, so we adopt the BN-based version, denoted as ResNet-18-BN, throughout our experiments.
Setting for ViT-Tiny
We construct a lightweight Vision Transformer model, ViT-Tiny, specifically tailored for federated learning on the CIFAR-100 dataset. The design is based on the standard ViT architecture (Dosovitskiy et al. 2020), with modifications to accommodate the small input size and limited data per client.
  • Input resolution: $32\times 32$
  • Patch size: $4\times 4$, resulting in 64 tokens per image
  • Embedding dimension: 192
  • Number of Transformer layers: 6
  • Number of attention heads: 3
  • Normalization: LayerNorm (applied before attention and MLP blocks)
  • Classification head: Linear projection to 100 classes (CIFAR-100)
  • Activation: GELU
  • Initialization: Xavier/Glorot for linear layers; sinusoidal positional encoding
To regularize training, we apply dropout (0.1) to both attention and MLP layers. All models are trained from scratch without pretraining.
Federated Learning Configuration.
We adopt the same federated learning setup as used in our ResNet experiments for a fair comparison:
  • Number of clients: 100
  • Client participation rate: 10%
  • Communication rounds: 300
  • Local update steps: 50 iterations per round
  • Local batch size: 50
Learning Rate Schedule.
We perform grid search to identify optimal learning rates for each algorithm:
  • FedAvg, SCAFFOLD, FedCM, and FedAdam: local learning rate of 0.1
  • Local Adam, FedLADA, Local AdamW, FedAdamW: local learning rate of 3e-4
  • FedAdam: global learning rate of 0.01
  • FedAdamW: global learning rate of 1.0
Weight Decay.
To ensure fair comparison under different regularization settings, we assign:
  • FedAvg, SCAFFOLD, FedCM, FedAdam: weight decay = 0.001
  • Local Adam, FedLADA, Local AdamW, FedAdamW: weight decay = 0.01
Optimizer.
We use Adam or AdamW as the local optimizer depending on the method. All optimizers use $\beta_{1}=0.9$, $\beta_{2}=0.999$, and weight decay of 0.01 when applicable.
Remarks.
Due to the smaller capacity of ViT-Tiny and limited data per client, we find that careful normalization (e.g., LayerNorm placement) and early learning rate warmup are beneficial. For future work, more advanced token-mixing techniques or hybrid CNN-ViT backbones may further improve performance in federated settings.
Swin Transformer Fine-tuning Settings
To demonstrate the effectiveness of our method on large-scale vision models, we conduct fine-tuning experiments using Swin Transformer-Tiny and ViT-Base on Tiny ImageNet and CIFAR-100. For both models, we initialize from official ImageNet-22K pre-trained weights (Liu et al. 2021; Dosovitskiy et al. 2020) to ensure consistency across methods.
Model Architecture: Swin-Tiny.
Swin-Tiny adopts a hierarchical Transformer structure that gradually reduces the spatial resolution while increasing the feature dimensions, mimicking a CNN-like pyramid:
  • Stage depth: [2, 2, 6, 2]
  • Number of attention heads: [3, 6, 12, 24]
  • Embedding dimensions: 96, 192, 384, 768 across stages
  • Patch size: $4\times 4$
  • Window size: 7
  • MLP ratio: 4
  • Normalization: LayerNorm
  • Positional encoding: Relative positional bias
  • Regularization: DropPath (with decay rate linearly scaled to depth)
We fine-tune all layers during federated training.
Data Preprocessing.
To align with the input resolution required by Swin and ViT, we resize images from both datasets to $224\times 224$ using bilinear interpolation. Standard data augmentation techniques such as random cropping, horizontal flipping, and RandAugment are applied locally at the client side.
Federated Learning Configuration.
To simulate a realistic cross-device setting, we configure:
  • Number of clients: 100
  • Client participation rate: 5% per communication round
  • Communication rounds: 100
  • Local update steps: 50 iterations
  • Batch size: 16
Learning Rate Configuration.
We apply grid search to find optimal learning rates and use cosine learning rate decay with no warmup unless otherwise stated:
  • FedAvg, SCAFFOLD, FedCM, FedAdam: local LR = 0.01
  • Local Adam, FedLADA, Local AdamW, FedAdamW: local LR = 1e-4
  • FedAdam: global LR = 0.01
  • FedAdamW: global LR = 1.0
Weight Decay.
To ensure fair comparison under different regularization settings, we assign:
  • FedAvg, SCAFFOLD, FedCM, FedAdam: weight decay = 0.001
  • Local Adam, FedLADA, Local AdamW, FedAdamW: weight decay = 0.01
Optimization.
Local optimizers are Adam or AdamW depending on the algorithm, with parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, and weight decay of 0.01. Cosine decay is applied to local learning rates over the 50 local steps per round. No learning rate warmup is used unless otherwise specified.
Remarks.
We find that Swin Transformer benefits from hierarchical attention and DropPath when training with limited local data. Our method shows stable convergence and avoids loss spikes often seen in large-scale federated fine-tuning. All models are implemented using the HuggingFace Transformers and Timm libraries.
RoBERTa-Base Fine-Tuning Settings
We fine-tune the RoBERTa-Base model using LoRA (Low-Rank Adaptation) on a subset of the GLUE benchmark. The LoRA adaptation is applied to the query and value projection matrices of the self-attention modules. The following table summarizes the hyperparameter settings used across tasks.
Table 9: Hyperparameter configuration for RoBERTa-Base with LoRA across GLUE tasks. LoRA is applied with $r=16$ to both query and value projections. All tasks use AdamW as the optimizer.
Method Setting MNLI SST-2 MRPC CoLA QNLI QQP RTE
RoBERTa-Base + LoRA Batch Size 32 32 32 32 32 32 32
Max Seq. Len. 128 128 128 128 128 128 128
LoRA Ranks $(r_{q}=r_{v})$ 16 16 16 16 16 16 16
LoRA Scaling $\alpha$ 32 32 32 32 32 32 32
Dropout / Weight Decay 0.1 / 0.01
Explanation. We use a uniform batch size of 32 and sequence length of 128 across all tasks. LoRA is configured with a rank of 16 and scaling factor $\alpha=32$. The optimizer is AdamW with a weight decay of 0.01 and dropout set to 0.1. No layer freezing is used; all LoRA-injected weights are trained, while the base RoBERTa backbone remains frozen.
Additional Federated Training Configuration
To evaluate our algorithm under a smaller-scale federation, we further conduct experiments with a reduced number of clients and adjusted participation parameters.
Federated Setup.
We simulate a federated learning environment with the following configuration:
  • Number of clients: 20
  • Client participation rate: 20% (i.e., 4 clients per round)
  • Communication rounds: 100
  • Local update steps: 50
  • Local batch size: 32
Learning Rate Schedule.
We apply grid search for local and global learning rates and use cosine learning rate decay across local updates:
  • FedAvg, SCAFFOLD, FedCM, FedAdam: local LR = 0.1
  • Local Adam, FedLADA, Local AdamW, FedAdamW: local LR = 1e-4
  • FedAdam: global LR = 0.01
  • FedAdamW: global LR = 1.0
Weight Decay.
To ensure fair comparison under different regularization settings, we assign:
  • FedAvg, SCAFFOLD, FedCM, FedAdam: weight decay = 0.001
  • Local Adam, FedLADA, Local AdamW, FedAdamW: weight decay = 0.01
Other Settings.
Adam and AdamW optimizers use $\beta_{1}=0.9$, $\beta_{2}=0.999$. Learning rates follow cosine decay without warmup.
Appendix D: Hessian Block Structure Partitioning
We now present the specific realization of FedAdamW on Transformers and other architectures. In those cases, we need to investigate the Hessian structure of these new architectures (like what we did for Transformers) and then develop the concrete partition algorithms.
Algorithm 3 Partition for Transformers
1:param_blocks = {}
2:for name, param in parameters do
3:  if ’embed’ or ’output’ in name then
4:   Partition param by tokens
5:   for i = 0...tokens-1 do
6:    param_blocks[name+i]=param[i]
7:   end for
8:  else if ’query’ or ’key’ in name then
9:   Partition param by heads
10:   for i = 0...heads-1 do
11:    param_blocks[name+i]=param[i]
12:   end for
13:  else if ’value’, ’attn.proj’, or ’mlp’ in name then
14:   Partition param by output neurons
15:   for i = 0...output_neurons-1 do
16:    param_blocks[name+i]=param[i]
17:   end for
18:  else
19:   param_blocks[name]=param
20:  end if
21:end for
22:return param_blocks
Algorithm 4 Partition for non-Transformers
1:param_blocks = {}
2:for name, param in parameters do
3:  param_blocks[name]=param
4:end for
5:return param_blocks
D.1 Motivation and Background
The motivation for using Hessian-based partitioning stems from the empirical observation that the Hessian matrix $\nabla^{2}f(x)$ of modern deep neural networks is often approximately block-diagonal, particularly in Transformer-like architectures. Each block corresponds to a structural or functional module (e.g., attention heads, feed-forward layers) which share similar curvature characteristics and learning dynamics.
Instead of transmitting the entire second-moment vector $v\in\mathbb{R}^{d}$, we propose to group parameters into $B$ blocks $\{v^{(1)},\ldots,v^{(B)}\}$ based on the underlying Hessian block structure, and communicate only the block-wise mean:
$\bar{v}^{(b)}=\text{mean}\left(v^{(b)}\right),\quad b=1,\ldots,B$
This strategy dramatically reduces communication overhead and improves robustness under data heterogeneity.
Figure 7 illustrates how we partition the model parameters in Adam-mini based on the Hessian structure of a Transformer, instead of relying on the default PyTorch parameter grouping. From subfigures (a)–(h), we observe that the Hessian matrices of different parameter groups exhibit a near block-diagonal structure, where each large block contains several smaller dense sub-blocks. Following Principle 1, we assign one distinct learning rate to each smallest dense sub-block in the Hessian. Specifically:
  • Query / Key: The Hessian shows a clear block-diagonal structure, with each block corresponding to an attention head. We therefore partition these parameters by head (e.g., $4$ heads $\rightarrow$ 44 blocks).
  • Value: The structure is less clear than for Query/Key, but hints at approximately as many diagonal blocks as the number of output neurons. By default, we partition by output neurons (e.g., $16$ neurons $\rightarrow$ $16$ blocks), though in some special cases it may be better to treat the entire value as a single block.
  • Attention projection (attn.proj) and MLP layers: The Hessian exhibits a block structure aligned with the output neurons. We thus partition by output neurons (e.g., $16$ or $32$ blocks depending on the layer).
  • Embedding and Output layers: The Hessian shows a near block-diagonal structure by token. We partition these parameters by token (e.g., $8$ tokens $\rightarrow$ 88 blocks).
This Hessian-based partitioning ensures that each dense Hessian sub-block is assigned at least one distinct learning rate, avoiding the overly coarse grouping of the PyTorch default partition.
(a) query (4 heads)
(b) key (4 heads)
(c) value (4 heads)
(d) attn.proj (16 neurons)
(e) mlp.fc_1 (32 neurons)
(f) mlp.proj (16 neurons)
(g) embed (8 tokens)
(h) output (8 tokens)
Figure 7: (a–h): The Hessian of different parameter blocks in a small Transformer at 1% training step. These sub-blocks show near block-diagonal structure with small dense matrices.
(a) Hessian of a MLP
(b) Hessian at initialization
(c) Hessian at 50% step
(d) Hessian at 100% step
Figure 8: The near-block-diagonal Hessian structure of neural nets in FL. (a) is the Hessian of an MLP reported in (Collobert 2004). (b–d) are the Hessians of a 1-hidden-layer MLP on CIFAR-100. The near-block-diagonal structure maintains throughout training in FL.
D.2 Practical Partitioning Strategy
  • Class 1: query and key. The Hessian of query and key have near-block-diagonal structures. The number of blocks equals the number of heads.
  • Class 2: attn.proj and MLPs. The Hessian of attn.proj and MLPs have block-diagonal structures. The number of blocks equals the number of output neurons.
  • Class 3: value. For value, the structure of the Hessian seems less clear. It appears to have 16 diagonal blocks (matching the number of output neurons), but the pattern is less obvious. This structure is significantly different from query and key, despite all using four heads. Additionally, the Hessian entries of value are about $10^{6}$ larger than those of query and key [^1] This might be one source of the heterogeneity of Hessian eigenvalues as reported by Zhang et al. (2024a). One possible reason is that value is positioned outside the softmax operator in the self-attention design, unlike query and key.
  • Class 4: embed and output. For these two layers, the Hessian sub-block has a near-block-diagonal structure, and the number of blocks equals the number of tokens.
We adopt a structural partitioning strategy aligned with model architecture. Concretely:
  • Transformer layers: Group by layer-wise modules. For instance, all weights in the self-attention block (e.g., $W_{Q}$, $W_{K}$, $W_{V}$), feed-forward networks ($W_{1}$, $W_{2}$), and layer norms are treated as separate blocks.
  • CNNs (e.g., ResNet): Group by convolutional layers or residual blocks.
  • Vision Transformers (ViT): Group patches and MLP heads independently.
Each block $v^{(b)}$ is a vector corresponding to a functional module. For stability, we use the following heuristic:
$B=\min\left(\texttt{max\_blocks},\texttt{floor}(d/\texttt{min\_block\_size})\right)$
where min_block_size is a hyperparameter (e.g., 512), and $d$ is the total parameter count.
D.3 Example: RoBERTa-Base Layer Partitioning
Consider a single Transformer layer in RoBERTa-Base:
  • self.attn.q_proj.weight   768 $\times$ 768
  • self.attn.k_proj.weight   768 $\times$ 768
  • self.attn.v_proj.weight   768 $\times$ 768
  • self.attn.out_proj.weight   768 $\times$ 768
  • self.fc1.weight, self.fc2.weight   3072 $\times$ 768
  • self.layernorm1.weight, self.layernorm2.weight   768
We assign each of the above to an independent block:
$\{v^{(1)}\}=\texttt{q\_proj},\quad\{v^{(2)}\}=\texttt{k\_proj},\quad\ldots,\quad\{v^{(7)}\}=\texttt{layernorm2}$
The server aggregates and broadcasts the scalar means $\{\bar{v}^{(1)},\ldots,\bar{v}^{(7)}\}$ for reuse in the next round. Despite this coarse granularity, our experiments show improved convergence and reduced variance.
Compare Theoretical Convergence Rate
Table 10: Comparison between Local AdamW, FedLADA and FedAdamW (ours).
Research work Local optimizer Convergence Rate Assumption Communication Cost
[…]
Table 10 compares Local AdamW, FedLADA, and our proposed FedAdamW in terms of the local optimizer, theoretical convergence rate, required assumptions, and communication cost.
Local AdamW adopts AdamW as the local optimizer but relies on the data heterogeneity assumption to guarantee convergence. Its theoretical convergence rate is
$\mathcal{O}\left(\sqrt{\frac{L\Delta(\sigma_{l}^{2}+\sigma_{g}^{2})}{SKR\epsilon^{2}}}+\frac{L\Delta}{R}\right),$
and the communication cost is $1\times$, indicating that only one model upload per communication round is required.
FedLADA uses Adam as the local optimizer, also depending on the data heterogeneity assumption, and shares the same convergence rate as Local AdamW. However, it requires additional communication for transmitting global update information, leading to a communication cost of $2\times$.
FedAdamW (ours) employs AdamW as the local optimizer and integrates global–local gradient alignment with second-moment mean aggregation to effectively mitigate the impact of data heterogeneity. Theoretically, it no longer requires the data heterogeneity assumption, and achieves an improved convergence rate:
$\mathcal{O}\left(\sqrt{\frac{L\Delta\sigma_{l}^{2}}{SKR\epsilon^{2}}}+\frac{L\Delta}{R}\right),$
while maintaining the same $1\times$ communication cost as Local AdamW.
In summary, FedAdamW relaxes the theoretical assumption requirements and improves the convergence rate bound without incurring extra communication cost, making it more suitable for federated learning scenarios with high data heterogeneity.
Table 11: Test accuracy, training loss on CIFAR-100 using ResNet-18 and ViT-Tiny over 300 communication rounds under Dir-0.6 and Dir-0.1 settings (100 clients, 10% participation, batch size 50, $K=50$).
Method ResNet-18 (Dir-0.6) ResNet-18 (Dir-0.1) ViT-Tiny (Dir-0.6) ViT-Tiny (Dir-0.1)
Test Acc Train Loss Test Acc Train Loss Test Acc Train Loss Test Acc Train Loss
Local AdamW $62.84_{\pm 0.08}$ 0.363 $58.97_{\pm 0.10}$ 0.794 $40.47_{\pm 0.09}$ 1.026 $36.86_{\pm 0.11}$ 1.954
FedAdamW(Algorithm 3) $65.86_{\pm 0.17}$ 0.116 $62.87_{\pm 0.15}$ 0.465 $42.23_{\pm 0.15}$ 0.385 $38.58_{\pm 0.17}$ 1.061
FedAdamW (Algorithm 2) 66.12±0.10 0.122 63.01±0.12 0.480 42.56±0.10 0.401 39.86±0.16 1.251
Compare FedAdamW (Algorithm 2) and FedAdamW (Algorithm 3)
Analysis of Table 11. Table 11 reports the test accuracy and final training loss on the CIFAR-100 dataset using two backbone architectures, ResNet-18 and ViT-Tiny, under two non-IID data partitions generated by the Dirichlet distribution with concentration parameters $\alpha=0.6$ and $\alpha=0.1$. The experiments are conducted over $300$ communication rounds with $100$ clients, $10\%$ client participation per round, a local batch size of $50$, and $K=50$ local optimization steps. The reported accuracy values are given as mean $\pm$ standard deviation over multiple runs.
(1) Effect of different optimization strategies. Local AdamW, which performs only local updates without parameter aggregation, yields the lowest accuracy across all settings, e.g., $58.97\%$ for ResNet-18 under $\alpha=0.1$. FedAdamW (Algorithm 3) significantly improves performance over Local AdamW in all cases, highlighting the benefit of federated aggregation. FedAdamW (Algorithm 2) achieves the best performance across all configurations, attaining both the highest test accuracy and the lowest training loss (e.g., $66.12\%$ accuracy and $0.122$ loss for ResNet-18 under $\alpha=0.6$).
(2) Impact of backbone architecture. ResNet-18 consistently outperforms ViT-Tiny across both Dirichlet settings, suggesting that convolutional architectures are better suited for CIFAR-100 in federated settings, especially under limited local data per client. Although ViT-Tiny exhibits lower absolute accuracy, the improvements from FedAdamW over Local AdamW remain consistent.
(3) Sensitivity to data heterogeneity. As the non-IID degree increases ($\alpha$ decreases from $0.6$ to $0.1$), performance drops for all methods, reflecting the challenge of heterogeneous data. However, FedAdamW (Algorithm 2) shows stronger robustness to heterogeneity, with smaller accuracy degradation compared to Local AdamW.
Conclusion. Overall, FedAdamW (Algorithm 2) delivers the best and most stable results across different architectures and data heterogeneity levels, demonstrating its effectiveness and generalizability in federated learning on CIFAR-100.