### v3

# FedAdamW: A Communication-Efficient Optimizer with Convergence and Generalization Guarantees for Federated Large Models

Abstract
AdamW has become one of the most effective optimizers for training large-scale models. We have also observed its effectiveness in the context of federated learning (FL). However, directly applying AdamW in federated learning settings poses significant challenges: (1) due to data heterogeneity, AdamW often yields high variance in the second-moment estimate $\boldsymbol{v}$; (2) the local overfitting of AdamW may cause client drift; and (3) reinitializing moment estimates ($\boldsymbol{v}$, $\boldsymbol{m}$) at each round slows down convergence. To address these challenges, we propose the first Federated AdamW algorithm, called FedAdamW, for training and fine-tuning various large models. FedAdamW aligns local updates with the global update using both a local correction mechanism and decoupled weight decay to mitigate local overfitting. FedAdamW efficiently aggregates the mean of the second-moment estimates to reduce their variance and reinitialize them. Theoretically, we prove that FedAdamW achieves a linear speedup convergence rate of $\mathcal{O}(\sqrt{(L\Delta\sigma_{l}^{2})/(SKR\epsilon^{2})}+(L\Delta)/R)$ without heterogeneity assumption, where $S$ is the number of participating clients per round, $K$ is the number of local iterations, and $R$ is the total number of communication rounds. We also employ PAC-Bayesian generalization analysis to explain the effectiveness of decoupled weight decay in local training. Empirically, we validate the effectiveness of FedAdamW on language and vision Transformer models. Compared to several baselines, FedAdamW significantly reduces communication rounds and improves test accuracy.
Codehttps://github.com/junkangLiu0/FedAdamW
Extended versionhttps://arxiv.org/pdf/2510.27486
Introduction
With the rapid growth of data and rising concerns over user privacy, traditional centralized training paradigms have become inadequate. Federated Learning (FL) (McMahan et al. 2017) offers a scalable and privacy-preserving framework that enables collaborative model training across decentralized clients without sharing raw data (Bian et al. 2025a; Wang et al. 2024; Bian et al. 2025b, 2024). As data becomes increasingly siloed, FL is a practical solution for large-scale distributed deep learning (Li et al. 2025a, 2023; An et al. 2022, 2024a, 2024b; Liu et al. 2023, 2025d, 2025e, 2025c).
However, recent trends in model design-particularly the rise of large-scale architectures such as GPT (Radford et al. 2018), RoBERTa (Liu et al. 2019), and Vision Transformers (ViT) (Dosovitskiy et al. 2020)—pose new challenges for existing FL algorithms. Specifically, the widely-used FedAvg algorithm, which relies on stochastic gradient descent (SGD(Bottou 2010) in local, struggles to efficiently train Transformer models. This is due to the slow convergence and poor adaptivity of SGD in Transformer models (Zhang et al. 2024c; Liu et al. 2025a), which have more complex architectures compared to CNNs. For example, components such as query, key, and value often require different learning rates to be trained effectively (Zhang et al. 2024c). In contrast, AdamW (Loshchilov et al. 2017), an adaptive optimizer with decoupled weight decay, has demonstrated superior performance in centralized training of large models based on Transformer (Vaswani et al. 2017; Liu et al. 2019), offering faster convergence and improved generalization, compared to Adam (Kingma and Ba 2014) and SGD.
Empirically, we also observe this advantage in FL: as shown in Figure 1, local training with AdamW (Local AdamW) converges significantly faster than Local SGD  (McMahan et al. 2017) for training various Transformer models . However, naively applying AdamW in FL leads to the following new challenges:
  • Challenge 1: High variance in second-moment estimate ($\boldsymbol{v}$). Due to non-i.i.d. data across clients, gradient noise leads to high variance in second-moment estimate.
  • Challenge 2: Local overfitting and client drift. While AdamW accelerates local training, it intensifies local overfitting. Under non-i.i.d. data, this manifests as client drift, severely hindering the global model’s performance.
  • Challenge 3: Moment estimate reinitialization. reinitializing the first- and second-moment estimates from scratch in every round hinders the convergence rate.
(a) ViT on CIFAR100
(b) GPT2 on Shakespeare
(c) BERT on CoLA
Figure 1: Performance of Local SGD and Local AdamW. For training ViT-Base, GPT2, and BERT (Liu et al. 2019), we carefully tune the learning rate. For training all these Transformer models, Local SGD is still significantly worse than Local AdamW.
These challenges motivate us to develop Federated AdamW (FedAdamW), a novel optimizer tailored for federated learning. FedAdamW addresses the above issues through two key designs: (1) a local correction mechanism that integrates global gradient estimates into the local update, effectively aligning local and global updates to reduce client drift; and (2) a moment aggregation strategy that aggregates the mean of second-moment estimates is theoretically grounded in the Hessian block structure, to reduce variance of $\boldsymbol{v}$ and avoid repeated initialization.
Our contributions are summarized as follows:
  • Empirical importance of AdamW and challenges in FL. We empirically demonstrate the effectiveness of the AdamW optimizer in federated settings, particularly for training Transformer models. Our analysis reveals three key challenges when applying AdamW in FL.
  • We propose FedAdamW, a principled FL algorithm tailored for adaptive optimizers. To address the above challenges, FedAdamW integrates global update estimate into local updates to mitigate overfitting and improve consistency. Inspired by the Hessian structure, we design a communication-efficient aggregation strategy that communicates the mean of second-moment across clients.
  • Theoretical guarantees with improved convergence and generalization. FedAdamW achieves a linear speedup convergence rate of $\mathcal{O}(\sqrt{(L\Delta\sigma_{l}^{2})/(SKR\epsilon^{2})}+(L\Delta)/R)$. To the best of our knowledge, this is the first federated adaptive optimization algorithm without requiring gradient heterogeneity assumption. Furthermore, we utilize the PAC-Bayesian theory to provide insights into the generalization benefits of decoupled weight decay and global-local alignment.
Related Work
$\bullet$ Heterogeneity Issues in Federated Learning. Data heterogeneity across clients is a fundamental challenge in FL. A range of algorithms have been proposed to mitigate the adverse effects of non-i.i.d. data distributions. For example, FedProx (Li et al. 2020) introduces a proximal term to restrict local updates; SCAFFOLD (Karimireddy et al. 2020) applies control variates to correct client drift; and FedCM (Xu et al. 2021) leverages client momentum to stabilize updates. FedNSAM (Liu et al. 2025b) analyzed the consistency between local and global flatness, FedBCGD  (Liu et al. 2024) proposed a communication-efficient accelerated block coordinate gradient method. FedSWA  (Liu et al. 2025a) further improved generalization under highly heterogeneous data via stochastic weight averaging.
$\bullet$ Adaptive Optimization in Centralized Settings. Adaptive gradient methods have demonstrated superior empirical performance over SGD in centralized settings, particularly for deep neural networks. Pioneering works include Adagrad (Duchi et al. 2011), Adam (Kingma and Ba 2014), AMSGrad (Reddi et al. 2019), and AdamW (Loshchilov et al. 2017). AdamW, in particular, decouples weight decay from gradient updates, offering improved generalization and training stability—attributes especially critical for Transformer models (Liu et al. 2019; Zhang et al. 2024c; Ouyang et al. 2025; Qian et al. 2024; Li et al. 2025b; Yang et al. 2025; Zhang et al. 2025, 2024b, 2024a; Wei et al. 2025; Zhou et al. 2023, 2024)
$\bullet$ Adaptive Optimization in Federated Learning. Recent efforts have explored integrating adaptive methods into FL. FedOpt (Reddi et al. 2020) incorporates server-side adaptivity using Adam and Yogi. FAFED (Wu et al. 2023) aggregates both the first- and second-moment estimates of Adam across clients to stabilize training. FedAMS  (Chen et al. 2020) shows that averaging the second-moment estimate of Adam is crucial to prevent divergence. More recently, Sun et al. (2023) proposed to only aggregate the second-moment estimate to reduce communication overhead. However, these works only conducted experiments on CNN models. These studies are all based on Adam, which performs poorly with large weight decay.
FL Problem Setup
FL aims to optimize model parameters with local clients, i.e., minimizing the following population risk:
$f(\boldsymbol{\boldsymbol{x}})=\frac{1}{N}\sum_{i=1}^{N}\left(f_{i}(\boldsymbol{\boldsymbol{x}}):=\mathbb{E}_{\xi_{i}\sim\mathcal{D}_{i}}\left[F_{i}\left(\boldsymbol{\boldsymbol{x}};\xi_{i}\right)\right]\right).$ (1)
The function $f_{i}$ represents the loss function on client $i$. $\mathbb{E}_{\xi_{i}\sim\mathcal{D}_{i}}[\cdot]$ denotes the conditional expectation with respect to the sample $\xi_{i}$. $\xi_{i}$ is drawn from distribution $\mathcal{D}_{i}$ in client $i$. NN is the number of clients.
Challenges of AdamW in FL
Despite the widespread use of AdamW (Loshchilov et al. 2017; Vaswani et al. 2017) in centralized deep learning, its adaptation to federated settings remains largely unexplored. In this section, we analyze three fundamental challenges that hinder its effectiveness in FL settings.
Challenge 1: High Variance in Second-Moment Estimates ($\boldsymbol{v}$). AdamW maintains a second-moment estimate ($\boldsymbol{v}$) to scale gradients adaptively, updated as:
$\boldsymbol{v}^{r,k}_{i}=\beta_{2}\boldsymbol{v}^{r,k-1}_{i}+(1-\beta_{2})\boldsymbol{g}^{r,k}_{i}\odot\boldsymbol{g}^{r,k}_{i},$ (2)
where $\boldsymbol{v}^{r,k}_{i}$ denotes the second-moment estimate maintained by client $i$ at local step $k$ of round $r$, $\boldsymbol{g}^{r,k}_{i}$ is the stochastic gradient, $\beta_{2}=0.999$ is the exponential decay rate for the second moment, and $\odot$ represents the element-wise (Hadamard) product in Algorithm 1. In FL, data heterogeneity leads to gradient heterogeneity. The squared stochastic gradients $\boldsymbol{g}^{r,k}_{i}\odot\boldsymbol{g}^{r,k}_{i}$ in Eq. (2) amplify the variance of $\boldsymbol{v}$ across clients in Figure 2 (a). This can cause instability and inefficient aggregation, especially when using non-i.i.d. data (Chen et al. 2020).
Challenge 2: Local Overfitting and Client Drift. While AdamW accelerates convergence through its adaptivity, it may exacerbate local overfitting. In FL, where each client minimizes its own local objective $f_{i}(\cdot)$, creating a natural gap between the local and global optima. Adaptive optimizers such as AdamW, with stronger update magnitudes, can drive clients further toward their local optima—diverging from the global direction. This leads to client drift as illustrated in Figure 2 (b), which manifests as inconsistencies in local models that degrade the global performance.
Challenge 3: Reinitialization Overhead. In FL, AdamW optimizer states are reinitialized from zero each round:
$\boldsymbol{m}^{r,0}_{i}\leftarrow\boldsymbol{0},\quad\boldsymbol{v}^{r,0}_{i}\leftarrow\boldsymbol{0}.$ (3)
Reinitializing moment estimates across rounds erases temporal memory, hindering the accumulation of adaptive statistics and slowing convergence, particularly in deep or large-scale models.
Algorithm 1 Local AdamW Algorithm
1:Initial model $\boldsymbol{x}^{0}$, $\beta_{1}=0.9,\beta_{2}=0.999,\epsilon=10^{-8}$, time step $t\leftarrow 0$, the number of all clients NN, each round selected clients $S$, weight decay $\lambda$.
2:for $r=1,\dots,R$ do
3:  for each selected client $i\in\{1,\dots,S\}$ in parallel do
4:   $\boldsymbol{x}_{i}^{r,0}\leftarrow\boldsymbol{x}^{r}$, $\boldsymbol{m}^{r,0}_{i}\leftarrow\boldsymbol{0}$, $\boldsymbol{v}^{r,0}_{i}\leftarrow\boldsymbol{0}$;
5:   for $k=1,\dots,K$ do
6:    $\boldsymbol{g}^{r,\tau}_{i}\leftarrow\nabla f_{i}(\boldsymbol{x}_{i}^{r,k};\xi_{i})$;
7:    $\boldsymbol{m}^{r,k}_{i}=\beta_{1}\boldsymbol{m}^{r,k-1}_{i}+\left(1-\beta_{1}\right)\boldsymbol{g}^{r,k}_{i}$;
8:    $\boldsymbol{v}^{r,k}_{i}=\beta_{2}\boldsymbol{v}^{r,k-1}_{i}+\left(1-\beta_{2}\right)\boldsymbol{g}^{r,k}_{i}\odot\boldsymbol{g}^{r,k}_{i}$;
9:    $\hat{\boldsymbol{m}}^{r,k}_{i}=\boldsymbol{m}^{r,k}_{i}/\left(1-\beta_{1}^{k}\right)$;
10:    $\hat{\boldsymbol{v}}^{r,k}_{i}=\boldsymbol{v}^{r,k}_{i}/\left(1-\beta_{2}^{k}\right)$;
11:    $\boldsymbol{x}_{i}^{r,k+1}\!=\!\boldsymbol{x}_{i}^{r,k}-\eta(\hat{\boldsymbol{m}}^{r,k}_{i}/(\sqrt{\hat{\boldsymbol{v}}^{r,k}_{i}}+\epsilon)-\lambda\boldsymbol{x}_{i}^{r,k})$;
12:   end for
13:   Communicate $(\boldsymbol{x}^{r,K}_{i}-\boldsymbol{x}^{r,0}_{i})$ to Server;
14:  end for
15:  $\boldsymbol{x}^{r+1}=\boldsymbol{x}^{r}+\frac{1}{S}\sum_{i=1}^{S}(\boldsymbol{x}^{r,K}_{i}-\boldsymbol{x}^{r,0}_{i})$;
16:  Communicate $(\boldsymbol{x}^{r+1})$ to Clients.
17:end for
Algorithm 2 FedAdamW Algorithm
1:Initial model $\boldsymbol{x}^{0}$, $\beta_{1}=0.9,\beta_{2}=0.999,\epsilon=10^{-8}$, time step $t\leftarrow 0$, the number of all clients NN, each round selected clients $S$, weight decay $\lambda$.
2:for $r=1,\dots,R$ do
3:  for each selected client $i\in\{1,\dots,S\}$ in parallel do
4:   $\boldsymbol{x}_{i}^{r,0}\leftarrow\boldsymbol{x}^{r}$, $\boldsymbol{m}^{r,0}_{i}\leftarrow\boldsymbol{0}$, $\boldsymbol{v}^{r,0}_{i}\leftarrow\boldsymbol{\bar{v}}^{r}$;
5:   for $k=1,\dots,K$ do
6:    $t\leftarrow t+1$;
7:    $\boldsymbol{g}^{r,k}_{i}\leftarrow\nabla f_{i}(\boldsymbol{x}_{i}^{r,k};\xi_{i})$;
8:    $\boldsymbol{m}^{r,k}_{i}=\beta_{1}\boldsymbol{m}^{r,k-1}_{i}+\left(1-\beta_{1}\right)\boldsymbol{g}^{r,k}_{i}$;
9:    $\boldsymbol{v}^{r,k}_{i}=\beta_{2}\boldsymbol{v}^{r,k-1}_{i}+\left(1-\beta_{2}\right)\boldsymbol{g}^{r,k}_{i}\odot\boldsymbol{g}^{r,k}_{i}$;
10:    Bias correction
11:    $\hat{\boldsymbol{m}}^{r,k}_{i}=\boldsymbol{m}^{r,k}_{i}/\left(1-\beta_{1}^{k}\right)$;
12:    $\hat{\boldsymbol{v}}^{r,k}_{i}=\boldsymbol{v}^{r,k}_{i}/\left(1-\beta_{2}^{t}\right)$;
13:    $\vartheta_{i}^{r,k}=1/(\sqrt{\hat{\boldsymbol{v}}_{i}^{r,k}}+\epsilon)$;
14:    Update model parameters $\boldsymbol{x}_{i}^{r,k+1}=\boldsymbol{x}_{i}^{r,k}-\eta(\hat{\boldsymbol{m}}^{r,k}_{i}\odot\vartheta_{i}^{r,k}+\alpha\boldsymbol{\Delta}_{G}^{r}-\lambda\boldsymbol{x}_{i}^{r,k})$;
15:   end for
16:   Communicate $(\boldsymbol{x}^{r,K}_{i}-\boldsymbol{x}^{r,0}_{i},\boldsymbol{\bar{v}}_{i}=\texttt{mean}\Big(\boldsymbol{v}^{r,K}_{i}\Big))$ to Server;
17:  end for
18:  $\boldsymbol{\Delta}_{G}^{r}=\frac{-1}{SK\eta}\sum_{i=1}^{S}(\boldsymbol{x}^{r,K}_{i}-\boldsymbol{x}^{r,0}_{i})$;
19:  $\boldsymbol{x}^{r+1}=\boldsymbol{x}^{r}+\frac{1}{S}\sum_{i=1}^{S}(\boldsymbol{x}^{r,K}_{i}-\boldsymbol{x}^{r,0}_{i})$;
20:  $\boldsymbol{\bar{v}}^{r+1}=\frac{1}{S}\sum_{i=1}^{S}\boldsymbol{\bar{v}}_{i}$;
21:  Communicate $(\boldsymbol{x}^{r+1},\boldsymbol{\bar{v}}^{r+1},\boldsymbol{\Delta}_{G}^{r})$ to Clients.
22:end for
(a) High Variance in $\boldsymbol{v}$
(b) Local AdamW client drift
Figure 2: Training on CIFAR-100 using ViT-Tiny. (a) Data heterogeneity causes high variance in second-moment estimates across clients of Local AdamW. (b) Local AdamW suffers from more severe client drift than Local SGD under non-i.i.d. data.
Our Algorithm: FedAdamW
(a) query (4 heads)
(b) key (4 heads)
(c) value (4 heads)
Figure 3: (a–c):Block-wise Hessian structure of Transformer parameters under FL. Visualizing the Hessian submatrices of query, key, and value heads. The near block-diagonal structure supports block-wise second-moment aggregation in FedAdamW.
Figure 4: Illustration of FedAdamW’s block-wise aggregation strategy Clients estimate local second-moment statistics and send block-wise means to the server, reducing communication cost.
Based on theoretical motivation, we propose an efficient improvement to AdamW called Federated AdamW (FedAdamW). To address Challenge 1, it was experimentally discovered that aggregating AdamW second-moment estimate can stabilize the training process (see Table 7 below). However, aggregating second-moment estimate leads to a double communication.
(Q1) How to efficiently aggregate $\boldsymbol{v}$?
We observe that the Hessian matrix in neural networks exhibits an approximate block-diagonal structure with several dense sub-blocks (Collobert 2004; Zhang et al. 2024d) as shown in Figure 3. In such a structure, a single learning rate can effectively capture the curvature within each block. Leveraging this, we propose a communication-efficient strategy that partitions the second-moment estimate $\boldsymbol{v}$ into $B$ blocks and transmits only the mean of each in Figure 4:
$\boldsymbol{\bar{v}}_{b}=\texttt{mean}(\boldsymbol{v}_{b}),\quad b=1,\dots,B.$ (4)
Block-wise Partitioning Strategy (ViT Example).
We group the parameters into semantically aligned classes that exhibit similar curvature patterns as shown in Figure 3:
  • Class 1: query and key. Query and Key parameters. Each block corresponds to one attention head.
  • Class 2: attn.proj and MLPs. Blocks align with output neurons in projection and feedforward layers.
  • Class 3: value. Structure is less regular but still shows diagonal blocks; curvature magnitude is notably higher (up to $10^{6}\times$), possibly due to its position after softmax.
  • Class 4: Embedding and output layers. Sub-blocks align with input tokens, forming near-diagonal Hessians.
CNNs (e.g., ResNet): Blocked by convolutional layers or residual blocks. This reduces the communication cost from billions of scalars to $B$ values while preserving adaptive behavior. Empirically, we find this approach also improves generalization in local optimization as shown in Table 7 below. See Appendix D for block partitioning details.
(Q2) How to overcome overfitting in Local AdamW?
To address local overfitting (i.e., Challenge 2), we adopt a stronger weight decay. Unlike Adam, AdamW employs decoupled weight decay, which improves generalization, particularly in federated settings (see Table 6 below). To further mitigate client drift under non-i.i.d. data, we incorporate a global update estimate into the local update rule:
$\boldsymbol{x}_{i}^{r,k+1}=\boldsymbol{x}_{i}^{r,k}-\eta\left(\hat{\boldsymbol{m}}^{r,k}_{i}\odot\vartheta_{i}^{r,k}-\lambda\boldsymbol{x}_{i}^{r,k}+\alpha\boldsymbol{\Delta}_{G}^{r}\right),$ (5)
where $\boldsymbol{\Delta}_{G}^{r}\!=\!\frac{-1}{SK\eta}\sum_{i=1}^{S}(\boldsymbol{x}_{i}^{r,K}\!-\!\boldsymbol{x}_{i}^{r,0})$ is the estimated global update. As shown in Figure 5, this alignment reduces the divergence of local models and improves global consistency.
(Q3) How to initialize second-moment estimates?
We find that initializing the second-moment estimate $\boldsymbol{v}$ with its aggregated mean $\boldsymbol{v}$ significantly accelerates convergence (see Table 7 below). In contrast, we reinitialize the first-moment estimate $\boldsymbol{m}$ to zero at each round. This is because $\boldsymbol{m}$ adapts quickly to recent gradients and does not require long-term accumulation to remain effective.
Figure 5: An illustration of local update in FedAdamW, which corrects client drift caused through global update guidance.
Theoretical Analysis
Convergence Analysis
In this part, we give the convergence theoretical analysis of our proposed FedAdamW algorithm. Firstly we state some standard assumptions for the non-convex function $f$.
Assumption 1 (Smoothness).
(Smoothness) The non-convex $f_{i}$ is a $L$-smooth function for all $i\in[m]$, i.e., $\|\nabla f_{i}(\boldsymbol{x})-\nabla f_{i}(\boldsymbol{y})\|\leq L\|\boldsymbol{x}-\boldsymbol{y}\|$, for all $\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{d}$.
Assumption 2 (Bounded Stochastic Gradient).
$\boldsymbol{g}_{i}^{r}=\nabla f_{i}(\boldsymbol{x}_{i}^{r},\xi_{i}^{r})$ computed by using a sampled mini-batch data $\xi_{i}^{r}$ in the local client $i$ is an unbiased estimator of $\nabla f_{i}$ with bounded variance, i.e., $\mathbb{E}_{\xi_{i}^{r}}[\boldsymbol{g}_{i}^{r}]=\nabla f_{i}(\boldsymbol{x}_{i}^{r})$ and $\mathbb{E}_{\xi_{i}^{r}}\|g_{i}^{r}-\nabla f_{i}(\boldsymbol{x}_{i}^{r})\|^{2}\leq\sigma_{l}^{2}$, for all $\boldsymbol{x}_{i}^{r}\in\mathbb{R}^{d}$.
Assumption 3 (Bounded Stochastic Gradient II).
Each element of stochastic gradient $\boldsymbol{g}_{i}^{r}$ is bounded, i.e., $\|\boldsymbol{g}_{i}^{r}\|_{\infty}=\|f_{i}(\boldsymbol{x}_{i}^{r},\xi_{i}^{r})\|_{\infty}\leq G_{g}$, for all $\boldsymbol{x}_{i}^{r}\in\mathbb{R}^{d}$ and any sampled mini-batch data $\xi_{i}^{r}$.
Assumption 4 (Bounded Heterogeneity).
The dissimilarity between local clients is bounded on the gradients, i.e., $\|\nabla f_{i}(\boldsymbol{x})-\nabla f(\boldsymbol{x})\|^{2}\leq\sigma_{g}^{2}$, for all $\boldsymbol{x}\in\mathbb{R}^{d}$.
These assumptions are standard in federated adaptive optimization literature (Fan et al. 2024; Sun et al. 2023).
Theorem 1 (Convergence for non-convex functions).
Under Assumptions 1, 2, and 3, if we take $g^{0}=0$$\beta_{1}=0,\lambda=0$ then FedAdamW converges as follows
$\frac{1}{R}\sum_{r=0}^{R-1}\mathbb{E}\left[\left\|\nabla f\left(\boldsymbol{x}^{r}\right)\right\|^{2}\right]\lesssim\mathcal{O}\left(\sqrt{\frac{L\Delta\sigma_{l}^{2}}{SKR\epsilon^{2}}}+\frac{L\Delta}{R}\right).$ (6)
Here $G_{0}:=\frac{1}{N}\sum_{i=1}^{N}\left\|\nabla f_{i}\left(\boldsymbol{x}^{0}\right)\right\|^{2}$$\Delta=f\left(\boldsymbol{x}^{0}\right)-f^{\star}$, $S$ is the number of participating clients per round, $K$ is the number of local iterations, and $R$ is the total number of communication rounds.
The proof is provided in Appendix A. The convergence rate of FedAdamW is faster than that of Local AdamW and FedLADA’s $\mathcal{O}\left(\sqrt{\frac{L\Delta(\sigma_{l}^{2}+\sigma_{g}^{2})}{SKR\epsilon^{2}}}+\frac{L\Delta}{R}\right)$, and we do not need Assumption 4. This is due to the suppression of local drift by the global update estimation $\boldsymbol{\Delta}_{G}^{r}$. We have verified this in Table 5 below.
Generalization Analysis
Theorem 2.
Assume the prior hypothesis $\boldsymbol{x}_{0}$ satisfies $\mathcal{P}_{\text{pre }}\sim\mathcal{N}(\mathbf{0},\rho\boldsymbol{I})$. Then the expected risk for the posterior hypothesis $\boldsymbol{x}\sim\mathcal{P}$ of FedAdamW learned on training dataset $\mathcal{D}_{tr}\sim\mathcal{D}$ with $n$ samples holds
$\mathbb{E}_{\boldsymbol{\xi}\sim\mathcal{D},\boldsymbol{x}\sim\mathcal{P}}[f(\boldsymbol{x},\boldsymbol{\xi})]-\mathbb{E}_{\boldsymbol{\xi}\in\mathcal{D}_{tr},\boldsymbol{x}\sim\mathcal{P}}[f(\boldsymbol{x},\boldsymbol{\xi})]\leq$
$\frac{\sqrt{8}}{\sqrt{n}}\left(\sum_{i=1}^{d}\log\frac{2\rho b\left(\sigma_{i}^{\frac{1}{2}}+\lambda\right)}{\eta}\!+\!\frac{\eta}{2\rho b}\sum_{i=1}^{d}\frac{1}{\sigma_{i}^{\frac{1}{2}}+\lambda}\!+\!c_{0}\!\right)^{\frac{1}{2}}$
, with at least probability $1-\tau$, where $\tau\in(0,1)$ and $c_{0}=\frac{1}{2\rho}\left\|\boldsymbol{x}_{*}\right\|^{2}-\frac{d}{2}+2\ln\left(\frac{2n}{\tau}\right)$. Here, $\sigma_{i}$ represents the local curvature (e.g., Hessian eigenvalue). $b$ is the batch size, $\eta$ is the learning rate, $\lambda$ is a weight decay parameter, $n$ is the training set size, and $d$ is the parameter dimension.
The proof is provided in Appendix B. Theorem 2 shows that the generalization error of FedAdamW can be upper bounded by $\mathcal{O}(1/\sqrt{n})$, where $n$ is the number of tatal data, consistent with classical results from PAC theory, stability, and uniform convergence (Shalev-Shwartz and Ben-David 2014). We further analyze the impact of the decoupled weight decay parameter $\lambda$ on this bound. As $\lambda$ increases, the first term $\sum_{i=1}^{d}\log 2\rho b(\sigma_{i}^{1/2}+\lambda)\eta^{-1}$ increases, while the second term $\frac{\eta}{2\rho b}\sum_{i=1}^{d}(\sigma_{i}^{1/2}+\lambda)^{-1}$ decreases. Although choosing the optimal $\lambda$ is challenging in practice, this trade-off suggests that tuning $\lambda$ appropriately can lead to a smaller generalization error, as shown in Table Table 6 below. This explains why FedAdamW often outperforms Local Adam (which corresponds to $\lambda=0$).
(a) ResNet18, Dir-0.6
(b) ResNet18, Dir-0.1
(c) ViT-Tiny, Dir-0.6
(d) ViT-Tiny, Dir-0.1
Figure 6: Training loss curves on CIFAR-100 using ResNet-18 and ViT-Tiny under Dir-0.1, Dir-0.6.
Method ResNet-18 (Dir-0.6) ResNet-18 (Dir-0.1) ViT-Tiny (Dir-0.6) ViT-Tiny (Dir-0.1) Comm
Test Acc Train Loss Test Acc Train Loss Test Acc Train Loss Test Acc Train Loss
FedAvg $64.08_{\pm 0.18}$ 0.376 $60.25_{\pm 0.20}$ 0.767 $32.36_{\pm 0.08}$ 2.350 $27.14_{\pm 0.12}$ 2.867 $1\times$
SCAFFOLD $65.01_{\pm 0.15}$ 0.365 $59.37_{\pm 0.16}$ 0.814 $32.17_{\pm 0.12}$ 2.295 $27.31_{\pm 0.11}$ 2.855 $2\times$
FedCM $48.69_{\pm 0.10}$ 1.305 $44.43_{\pm 0.08}$ 1.645 $26.33_{\pm 0.06}$ 2.681 $23.18_{\pm 0.15}$ 3.038 $1\times$
Local Adam $60.98_{\pm 0.28}$ 1.598 $58.88_{\pm 0.20}$ 0.975 $38.69_{\pm 0.16}$ 2.082 $29.88_{\pm 0.08}$ 2.961 $1\times$
FedAdam $63.77_{\pm 0.13}$ 0.562 $61.62_{\pm 0.16}$ 0.707 $28.77_{\pm 0.12}$ 2.709 $23.49_{\pm 0.15}$ 3.084 $1\times$
FedLADA $65.07_{\pm 0.18}$ 0.671 $59.93_{\pm 0.21}$ 0.556 $37.31_{\pm 0.16}$ 2.127 $35.33_{\pm 0.16}$ 2.678 $2\times$
Local AdamW $62.84_{\pm 0.08}$ 0.363 $58.97_{\pm 0.10}$ 0.794 $40.47_{\pm 0.09}$ 1.026 $36.86_{\pm 0.11}$ 1.954 $1\times$
FedAdamW 66.12±0.10 0.122 63.01±0.12 0.480 42.56±0.10 0.401 39.86±0.16 1.251 $1\times$
Table 1: Test accuracy, training loss, and communication cost of each method on CIFAR-100 using ResNet-18 and ViT-Tiny over 300 communication rounds under Dir-0.6 and Dir-0.1 settings (100 clients, 10% participation, batch size 50, $K=50$).
Experiments
Datasets. We evaluate FedAdamW on both vision and language tasks. (i) For image classification, we use CIFAR-100 (Krizhevsky et al. 2009), and Tiny ImageNet (Le and Yang 2015). (ii) For NLP tasks, we adopt benchmark datasets from the GLUE benchmark, including SST-2, QQP. To simulate data heterogeneity across clients, we follow the Dirichlet partitioning scheme (Hsu et al. 2019), where a Dir-0.6 corresponds to a low heterogeneity and Dir-0.1 implies high heterogeneity.
Model Architectures. We explore a variety of model types: (i) ResNet-18 (He et al. 2016) as a representative convolutional neural network (CNN), (ii) Swin Transformer (Liu et al. 2021) and ViT-Tiny (Dosovitskiy et al. 2020) for Vision Transformers, and (iii) RoBERTa-Base (Liu et al. 2019) for large-scale language model.
Baselines. We compare our method against state-of-the-art FL algorithms: FedAvg (McMahan et al. 2017), SCAFFOLD (Karimireddy et al. 2020), FedCM (Xu et al. 2021), FedAdam (Reddi et al. 2020), FedLADA (Sun et al. 2023), Local Adam and Local AdamW.
Hyperparameter Settings. For FedAvg, SCAFFOLD, FedCM, FedAdam, the $lr$ is selected from $\{10^{-2},\ 3\times 10^{-2},\ 5\times 10^{-2},\ 10^{-1},\ 3\times 10^{-1}\}$, with a weight decay of $0.001$. For FedAdamW, FedLADA, Local Adam and Local AdamW, the $lr$ is selected from $\{10^{-4},\ 3\times 10^{-4},\ 5\times 10^{-4},\ 8\times 10^{-4},\ 10^{-3}\}$, with weight decay $0.01$ or $0.001$, $\beta_{1}=0.9$, $\beta_{2}=0.999$. We apply cosine learning rate decay, and set FedAdamW to $\boldsymbol{\alpha}\!=\!0.5$, weight decay $\boldsymbol{\lambda}\!=\!0.01$. Additional hyperparameter configurations are detailed in Appendix C. We release all code, configuration files to ensure full reproducibility. All results are averaged over 5 runs with std reported.
Questions. Our experiments are designed to answer the following: Q1. Does Local AdamW outperform Local SGD when training Transformer models? Q2. Can FedAdamW effectively address the three challenges identified for AdamW in FL? Q3. Is FedAdamW generally effective across both CNNs and Transformers? Q4. Are individual components of FedAdamW—such as global update correction, decoupled weight decay, and block-wise $\boldsymbol{v}$ averaging—empirically beneficial? Q5. Do our theoretical findings (Theorems 1 and 2) align with empirical results?
Method CIFAR-100 Tiny ImageNet
Test Acc Train Loss Test Acc Train Loss
FedAvg $80.02_{\pm 0.28}$ 0.588 $80.38_{\pm 0.22}$ 0.826
SCAFFOLD $81.30_{\pm 0.18}$ 0.514 $82.41_{\pm 0.18}$ 0.650
FedCM $82.38_{\pm 0.19}$ 0.565 $83.18_{\pm 0.19}$ 0.522
Local Adam $79.75_{\pm 0.26}$ 0.534 $73.63_{\pm 0.28}$ 1.045
FedAdam $77.48_{\pm 0.19}$ 0.651 $78.20_{\pm 0.22}$ 0.834
FedLADA $74.64_{\pm 0.18}$ 0.598 $70.95_{\pm 0.19}$ 0.944
Local AdamW $83.35_{\pm 0.10}$ 0.381 $80.26_{\pm 0.12}$ 0.686
FedAdamW 85.85±0.08 0.285 85.23±0.10 0.446
Table 2: Comparison of test accuracy and training loss for Swin Transformer under Dir-0.1 with 100 communication rounds(100 clients, 5% participation, batch size 16, K=50K\!=\!50).
Method (Dir-0.8) CoLA RTE SST-2 QQP MRPC QNLI MNLI Avg Acc.
FedAvg 56.12±0.18 48.72±0.25 93.66±0.10 85.87±0.14 86.00±0.12 90.21±0.09 83.22±0.17 77.68±0.17
SCAFFOLD 57.79±0.21 51.62±0.28 93.15±0.11 84.25±0.15 86.11±0.13 90.32±0.10 83.49±0.18 77.82±0.17
FedCM 56.29±0.16 64.98±0.22 93.25±0.12 83.19±0.17 85.56±0.13 88.13±0.15 78.90±0.19 78.33±0.18
Local Adam 56.08±0.19 62.81±0.23 93.80±0.09 85.07±0.13 84.55±0.14 88.57±0.12 82.62±0.16 79.07±0.15
FedAdam 55.26±0.20 58.12±0.26 93.26±0.10 85.12±0.13 86.11±0.11 89.21±0.09 83.16±0.17 78.32±0.16
FedLADA 50.00±0.24 57.40±0.25 93.57±0.11 85.88±0.14 82.59±0.15 89.76±0.10 82.99±0.16 77.17±0.17
Local AdamW 56.45±0.17 54.15±0.27 93.57±0.09 86.93±0.12 86.27±0.11 90.73±0.08 84.26±0.14 78.91±0.15
FedAdamW (ours) 58.21±0.15 66.48±0.20 94.03±0.08 87.62±0.11 86.76±0.10 90.88±0.07 84.55±0.13 81.79±0.14
Table 3: Test accuracy (%) using RoBERTa-Base with LoRA across seven GLUE tasks over 100 communication rounds.
Results on Convolutional Neural Networks
Training on CIFAR-100 with ResNet-18.
Table 1 and Figure 6 present the test accuracy and training loss on CIFAR-100 using ResNet-18. FedAdamW achieves the best performance under both Dir-0.6 and Dir-0.1 settings, reaching a top accuracy of 66.12% and 63.01%, respectively. It also attains the lowest training loss (0.122 and 0.480), demonstrating faster and more stable convergence. Compared to other adaptive baselines such as FedAdam, FedAdamW shows superior generalization under data heterogeneity, confirming its effectiveness in CNNs (Q3).
Results on Transformer Models
Training on CIFAR-100 with ViT-Tiny. Table 1 and Figure 6 show FedAdamW achieves the best performance across both heterogeneity levels, with test accuracies of 42.56% (Dir-0.6) and 38.25% (Dir-0.1), and the lowest training loss (0.401 and 1.251), confirming its efficient convergence (Q5). Compared to Local AdamW, it provides consistent improvements in both accuracy and stability (Q1,Q2, Q3). Moreover, other adaptive baselines such as FedAdam and FedLADA perform significantly worse under high heterogeneity, highlighting the effectiveness of global update correction and decoupled weight decay (Q4). These results validate that FedAdamW is particularly effective for federated vision Transformers under non-i.i.d. conditions. The small dataset CIFAR100 is difficult to support the performance of ViT, resulting in lower accuracy. Therefore, we continued to test on the pretrained model.
Fine-tuning Results on Swin Transformer. Table 2 reports results on Swin Transformer under Dir-0.1. FedAdamW achieves the highest test accuracy on both CIFAR-100 (85.85%) and Tiny ImageNet (85.23%), while also attaining the lowest training loss, reflecting faster convergence and improved generalization. Compared to other adaptive baselines such as FedAdam and FedLADA, FedAdamW consistently outperforms across both datasets, demonstrating its effectiveness in fine-tuning large Transformer models under non-i.i.d. conditions.
Fine-tuning Results on LLMs. Table 3 summarizes results on the GLUE benchmark using RoBERTa-Base with LoRA, 20 clients, 20% participation, batch size 32, $K=50$, rank==16. FedAdamW achieves the highest average accuracy of 81.79%, outperforming strong baselines such as FedAvg (77.68%) and Local AdamW (78.91%). It is particularly strong on challenging tasks like RTE and QQP, exceeding the next best methods by +1.50% and +1.74%, respectively.
Ablation Study
Impact of A1, A2, A3. Table 4 summarizes the effect of removing key components in FedAdamW. We draw the following observations: $\bullet$ Removing second-moment aggregation (A1) significantly degrades performance, indicating that mean($\boldsymbol{v}$) aggregation stabilizes adaptive updates across clients. $\bullet$ Without global gradient alignment (A2), the local models drift apart, resulting in higher train loss and lower generalization. $\bullet$ The use of standard (non-decoupled) weight decay (A3) leads to suboptimal regularization, confirming the necessity of decoupled weight decay for Transformer-based federated training. $\bullet$ The complete FedAdamW (A4) consistently outperforms all ablated versions, validating the effectiveness of our joint design.
Impact of $\alpha$. Table 5 evaluates the effect of the global update alignment parameter $\alpha$ in FedAdamW. As predicted by our convergence analysis (Theorem 1), incorporating global update direction helps suppress client drift and accelerates convergence. We observe that $\alpha=0.5$ yields the best performance, striking a balance between local adaptivity and global consistency, in line with our theoretical insight (Q5).
Variant Test Acc (%) Train Loss
A1: w/o $\boldsymbol{\bar{v}}$ (no moment agg.) $37.51_{\pm 0.12}$ 1.504
A2: w/o $\boldsymbol{\Delta}_{G}$ (no global align.) $37.42_{\pm 0.14}$ 1.621
A3: w/o decoupled weight decay $38.25_{\pm 0.16}$ 1.356
A4: FedAdamW (Full) $\mathbf{39.86}_{\pm 0.16}$ $\mathbf{1.251}$
Table 4: Ablation study of on CIFAR-100 using ViT-Tiny (Dir-0.1, 300 rounds).
$\alpha$ (Dir-0.1) 0.00 0.25 0.50 0.75 1.00
Test Acc (%) 36.86 37.93 39.86 37.47 36.25
Train Loss 1.954 1.586 1.251 1.362 1.491
Table 5: Impact of $\alpha$ using ViT-Tiny on CIFAR-100.
Impact of weight decay $\lambda$. Table 6 shows that decoupled weight decay, as used in AdamW and FedAdamW, consistently improves test accuracy over standard Adam. FedAdamW generalizes well across all $\lambda$ values, with $\lambda=0.01$ performing best. This aligns with our PAC-Bayesian analysis (Theorem 2), where an appropriate $\lambda$ balances regularization and curvature for better generalization (Q5).
$\lambda$ 0.0005 0.001 0.005 0.010 0.020
Local Adam 28.86 29.88 18.65 8.56 4.05
Local AdamW 35.82 36.12 36.54 36.86 36.28
FedAdamW 38.26 39.24 39.55 39.86 38.56
Table 6: Ablation on weight decay $\lambda$ using ViT-Tiny on CIFAR-100 (Dir-0.1). FedAdamW consistently outperforms local baselines across a range of $\lambda$ values.
Impact of Aggregation Strategy. Table 7 shows that our strategy, Agg-mean-$\boldsymbol{v}$, achieves the best balance between accuracy and communication cost. While Agg-$\boldsymbol{v}$ improves performance by reducing variance, full aggregation (Agg-$\boldsymbol{vm}$) introduces excessive communication with marginal gains. In contrast, Agg-mean-$\boldsymbol{v}$ attains similar benefits with only $\mathcal{O}(B)$ communication, where $B$ is the number of blocks, demonstrating its scalability and effectiveness in stabilizing updates.
Aggregation Strategy Acc Train Loss Comm(↑)
NoAgg 36.86±0.11 1.954 5.7M
Agg-$\boldsymbol{m}$ 37.12±0.13 1.854 11.4M
Agg-$\boldsymbol{v}$ 38.01±0.12 1.652 11.4M
Agg-$\boldsymbol{vm}$ (FullAgg) 38.12±0.12 1.645 17.1M
Agg-mean-$\boldsymbol{v}$ 38.15±0.10 1.601 5.7M
Table 7: Ablation study of moment aggregation strategies of Local AdamW on CIFAR-100 with ViT-Tiny under Dir-0.1.
Conclusion
In this work, we proposed a novel federated optimization algorithm (FedAdamW) for training large-scale Transformer models. FedAdamW tackles the key challenges of applying AdamW in federated settings, including high variance in second-moment estimates, local overfitting under non-i.i.d. data, and inefficiencies from frequent reinitialization. It integrates second-moment aggregation, global update correction, and decoupled weight decay. We provided convergence analysis under non-convex and use the PAC Bayesian theory to support its generalization benefits. Extensive experiments on vision and language tasks verified that FedAdamW consistently outperforms strong FL baselines, especially on Transformer architectures, demonstrating its practical and theoretical strengths. We believe FedAdamW opens a new direction for adapting modern optimizers to FL such as LAMB (Chen et al. 2023) or Lion (Chen et al. 2023).
Acknowledgments
This work was supported by the National Natural Science Foundation of China (Nos. 62276182, 62276004), Peng Cheng Lab Program (No. PCL2023A08), Tianjin Natural Science Foundation (Nos. 24JCYBJC01230, 24JCYBJC01460), Tianjin Municipal Education Commission Research Plan (No. 2024ZX008).